### v1

# Kolmogorov-Arnold Attention: Is Learnable Attention Better For Vision Transformers?

Abstract
Kolmogorov-Arnold networks (KANs) are a remarkable innovation consisting of learnable activation functions with the potential to capture more complex relationships from data. Although KANs are useful in finding symbolic representations and continual learning of one-dimensional functions, their effectiveness in diverse machine learning (ML) tasks, such as vision, remains questionable. Presently, KANs are deployed by replacing multilayer perceptrons (MLPs) in deep network architectures, including advanced architectures such as vision Transformers (ViTs). In this paper, we are the first to design a general learnable Kolmogorov-Arnold Attention (KArAt) for vanilla ViTs that can operate on any choice of basis. However, the computing and memory costs of training them motivated us to propose a more modular version, and we designed particular learnable attention, called Fourier-KArAt. Fourier-KArAt and its variants either outperform their ViT counterparts or show comparable performance on CIFAR-10, CIFAR-100, and ImageNet-1K datasets. We dissect these architectures’ performance and generalization capacity by analyzing their loss landscapes, weight distributions, optimizer path, attention visualization, and spectral behavior, and contrast them with vanilla ViTs. The goal of this paper is not to produce parameter- and compute-efficient attention, but to encourage the community to explore KANs in conjunction with more advanced architectures that require a careful understanding of learnable activations. Our open-source code and implementation details are available on: subhajitmaity.me/KArAt
Multi-Head Self-Attention, Vision Transformers, Kolmogorov-Arnold Network, Kolmogorov-Arnold Transformers, Kolmogorov-Arnold Attention
Figure 1: Model parameters vs. Top-1 Test accuracy in ImegeNet-1K training of vanilla ViTs (Dosovitskiy et al., 2020), Vision KAN (DeiT+KAN) by (Chen et al., 2024), ViT+KAN and Kolmogorov-Arnold Transformer (KAT) by (Yang & Wang, 2025).
1 Introduction
Artificial general intelligence has become a rapidly growing research direction, and Kolmogorov-Arnold Network (KAN) (Liu et al., 2024) marks a remarkable innovation to that. KANs with learnable activation functions can potentially capture more complex relationships and facilitate meaningful interaction between the model and human intuition. After training a KAN on a specific problem, researchers can extract the learned univariate functions the model uses to approximate complex multivariable functions. By studying these learned functions, researchers can gain insights into the underlying relationships from the data and refine the model. KANs exhibit state-of-the-art performance in finding symbolic function representations (Yu et al., 2024), continual learning of one-dimensional functions (Liu et al., 2025).
KANs were integrated with neural network architectures, the primary of them being conventional MLPs or convolution neural networks (CNNs). E.g., (Ferdaus et al., 2024; Bodner et al., 2024; Drokin, 2024; Abueidda et al., 2025; Wang et al., 2025) combine KAN with CNNs, (Li et al., 2024) combine KAN with U-Net, etc. Interestingly, (Yu et al., 2024) claimed to make the first fairer comparison between KANs and MLPs on multiple ML tasks on small-scale datasets. By setting the same parameter counts in both networks, (Yu et al., 2024) concluded that KANs with B-Spline basis significantly underperform compared to their MLP counterparts in all ML tasks except symbolic formula representation; a similar claim reinstated in (Azam & Akhtar, 2024) for vision. But KAN’s exploration of more advanced architectures such as Transformers, remains limited. (Chen et al., 2024) replace the MLP layers inside the encoder blocks of data-efficient image Transformers (DeiTs) (Touvron et al., 2021) with KANs and proposed DeiT+KAN, (Yang & Wang, 2025) proposed two variants; ViT+KAN that replaces the MLP layers inside ViT’s encoder blocks, and Kolmogorov-Arnold Transformer (KAT), albeit similar to ViT+KAN but with a refined group-KAN strategy.
While KANs were claimed to be not superior to simple MLPs in vision tasks (Yu et al., 2024; Azam & Akhtar, 2024), we were curious to find out how ViTs can see images through the lens of KANs. In that effort, Figure 1 demonstrates the performance of vanilla ViTs (ViT tiny, ViT small (Wu et al., 2022; Steiner et al., 2022), and ViT base (Dosovitskiy et al., 2020)) and their KAN counterparts on ImageNet-1K (Deng et al., 2009) training. While different variants of DeiT+KAN (Chen et al., 2024) and ViT+KAN (Yang & Wang, 2025) show a 5%–18% drop in the Top-1 accuracy compared to vanilla ViTs and DeiTs in ImageNet-1K, the KATs (Yang & Wang, 2025) show approximately 1%–2.5% gain in Top-1 accuracy compared to vanilla ViTs and DeiTs by keeping about the same number of parameters. These results show KANs require further investigation in more advanced architectures such as Transformers. Nevertheless, first, we want to understand why there is a discrepancy in these two seemingly equivalent implementations.
In search for an answer, we realized that while DeiT+KAN and ViT+KAN replace the MLP layer with KAN in the Transformer’s encoder block, KAT implements a sophisticated group-KAN strategy that reuses a learnable function among a group of units in the same layer and chooses different bases for different encoder blocks. This strategy helps improve the accuracy by keeping almost similar parameter counts. Therefore, simply replacing MLPs with KANs might not guarantee better generalizability, but a properly designed KAN could. This perspective on the KANs’ performance in conjunction with Transformers opens many research questions. Importantly, we also note that the community remains put from designing a learnable multihead attention module, the heart of the Transformers. Therefore, in the rise of second-generation Transformers, such as Google’s TITAN (Behrouz et al., 2024) and SAKANA AI’s Transformer2 (Sun et al., 2025) that mimic the human brain, we want to ask: Is it worth deploying learnable multi-head self-attention to the (vision) Transformers?
Figure 2: (a) The standard softmax self-attention for $i^{\rm th}$ head in the $j^{\rm th}$ encoder block. (b) The Kolmogorov-Arnold Attention (KArAt) replaces the softmax with a learnable function $\Phi^{i,j}$. (c) The ideal KArAt is with an operator matrix, $\Phi^{i,j}$ with $N\times N$ learnable units that act on each row of ${\cal A}^{i,j}$. (d) While $\Phi^{i,j}\in\mathbb{R}^{N\times N}$ is impossible to implement due to computational constraints, our architecture uses an operator $\widehat{\Phi}^{i,j}$ with $N\times r$ learnable units $r\ll N$, followed by a linear projector with learnable weights $W\in\mathbb{R}^{r\times N}$.
Finally, we note that there is a line of research that proposes efficient attention mechanisms in terms of sparse attention (Rahimian et al., 2024; Yun et al., 2020; Shi et al., 2021; Kovaleva et al., 2019; Zhang et al., 2021; Zaheer et al., 2020; Guo et al., 2019) or linear/kernelized attention (Han et al., 2025; Nguyen et al., 2023, 2021, 2022b; Lu et al., 2021; Nguyen et al., 2022a), to remedy computational and memory complexities of attention calculation. In contrast, this paper tries to understand how learnable multi-head self-attention (MHSA) modules can perform over regular self-attention used in ViTs, a defining technology in vision in the last six years. We list our contributions as follows:
Designing a Learnable Attention Module for ViTs. Since language is significantly more complicated, less structured, and has a higher information entropy than an image, we limit this pilot study to the ViTs (Dosovitskiy et al., 2020) for image classification. In §3, we propose a general learnable Kolmogorov-Arnold multi-head self-attention mechanism or KArAt for vanilla ViTs that can operate on any basis. However, the computing and memory costs of training them motivated us to propose a more modular version, and we designed a particular MHSA mechanism using Fourier basis, called Fourier KArAt and its variants; see §4 for details.
Benchmarking, Evaluation, and Analysis. In §5, we benchmark different variants of Fourier KArAt on CIFAR-10, CIFAR-100, (Krizhevsky et al., 2009) and ImageNet-1K (Deng et al., 2009) datasets and evaluate their performance against vanilla ViTs—ViT tiny, ViT small (Wu et al., 2022; Steiner et al., 2022), and ViT base (Dosovitskiy et al., 2020). We dissect these architectures’ performance and generalization capacity by analyzing the loss landscapes, weight distributions, optimizer path, attention visualization, and spectral behavior and contrast them with vanilla ViTs.
2 Background
Multi-layer Perceptrons (MLPs) consist of $L$ layers, where the $l^{\rm th}$-layer is defined for input $x^{[l-1]}\in\mathbb{R}^{d_{l-1}}$ as
${a^{[l]}=\Phi^{[l]}(W_{l}x^{[l-1]}+b_{l}),\ W_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}},b_{l}\in\mathbb{R}^{d_{l}}.}$
By convention, $a^{[0]}=x^{[0]}$ denotes the input data. In MLP, the nonlinear activation functions, $\Phi(\cdot)$ are fixed (Bengio et al., 2017). For supervised tasks, given a training dataset $D$ with $N$ elements of the form (input, ground-truth) pairs, $\{(x_{i},y^{\star}_{i})\}_{i=1}^{N}$, the loss, $\mathcal{L}(X,Y^{\star}|\mathcal{W})=d_{Y}(a^{[L]},Y^{\star})$ (metric induced by the space $\mathbb{R}^{d_{L}}$) is calculated in the forward pass. The layerwise weights, $\{(W_{l},b_{l})\}_{l\in[L]}$ are learned by minimizing $\mathcal{L}(X,Y^{\star}|\mathcal{W})$.
Kolmogorov-Arnold Network (KAN) (Liu et al., 2025) is a neural network involving learnable activations parametrized by a chosen set of basis functions defined on a set of grid points or knots. The idea for this network stems from the Kolmogorov-Arnold Representation Theorem.
Theorem 1 (Kolmogorov-Arnold Representation Theorem).
(Kolmogorov, 1956) For any multivariate continuous function, $f:[0,1]^{n}\to\mathbb{R}$, there exists a finite composition of continuous single-variable functions, $\phi_{q,p}:[0,1]\to\mathbb{R},\Phi_{q}:\mathbb{R}\to\mathbb{R}$ such that
$f(x)=f(x_{1},x_{2},\dots,x_{n})=\sum\limits_{q=1}^{2n+1}\Phi_{q}\bigg{(}\sum\limits_{p=1}^{n}\phi_{q,p}(x_{p})\bigg{)}.$
Theorem 1 describes an exact finite expression using 2 layers, but, this can be extended to a multi-layer KAN; it is analogous to the Universal Approximation Theorem (Hornik et al., 1989).
For input $x^{[0]}$, an $L$-layer KAN is given as $\text{KAN}(x^{[0]})=\Phi^{L}\circ\cdots\circ\Phi^{l}\circ\cdots\circ\Phi^{1}(x^{[0]})$. In each KAN layer, the activation function $\Phi$ is learnable, and ${\Phi^{[l]}=[\Phi_{ij}^{[l]}]}$ operates on ${x^{[l-1]}\in\mathbb{R}^{d_{l-1}}}$ to produce ${x^{[l]}\in\mathbb{R}^{d_{l}}},$ such that:
${x^{[l]}_{i}=\Phi^{[l]}_{i:}(x^{[l-1]})=\sum\limits_{j=1}^{d_{l-1}}\Phi^{[l]}_{ij}(x^{[l-1]}_{j})}.$
Originally, (Liu et al., 2025) chose the B-spline basis functions. That is, the activation functions were defined to be $\phi(x)=w(b(x)+\text{spline}(x))$, where $b(x)=\text{silu}(x)=\frac{x}{1+e^{-x}}$ and $\text{spline}(x)=\sum\limits_{i}c_{i}B_{i}(x)$, with $B_{i}(\cdot)$ being one of the $k$-th degree B-spline basis functions parametrized by the $G$ grid points on a uniform grid, $[-I,I]$. In this case, the representation for each function is given as:
$\Phi_{ij}^{[l]}(x^{[l-1]}_{j})=w^{[l]}_{ij}(\text{silu}(x^{[l-1]}_{j})+\sum\limits_{m=1}^{G+k-1}c^{[l]}_{ijm}B_{ijm}(x^{[l-1]}_{j})).$
The weights $\{([c^{[l]}_{ijm}],[w^{[l]}_{ij}])\}_{l=1}^{L}$ are learned layerwise, similar to MLPs, by minimizing by final loss for a supervised learning task. However, the bases for the activations could be Fourier (Mehrabian et al., 2024; Xu et al., 2024), wavelet basis functions (Bozorgasl & Chen, 2024), fractals (Yang & Wang, 2025) etc.
Multi-head self-attention (MHSA) in ViTs (Dosovitskiy et al., 2020). The encoder-only vanilla ViT architecture is inspired by the Transformer proposed in (Vaswani et al., 2017). For simplicity, we do not mention the details of the layer normalization and other technicalities. Our central focus is the MHSA architecture.
Let ${X\in\mathbb{R}^{H\times W\times C}}$ be an input image of resolution ${H\times W}$. ViT uses an $p\times p$ patch and generates $N=\tfrac{HW}{p^{2}}$ input tokens. Each token is vectorized along $C$ channels to produce a matrix ${X_{T}\in\mathbb{R}^{N\times p^{2}C}}$, where each row, $X_{T_{i,:}}$, represents a flattened patch across $C$ channels. A learnable embedding matrix, $E\in\mathbb{R}^{p^{2}C\times d}$ projects each row ${X_{T}}_{i,:}$ to an embedding vector in $\mathbb{R}^{d}$ such that After appending the ground-truth class, $x_{{\cal G}}$ and adding the positional encoding matrix, $P_{E}\in\mathbb{R}^{(N+1)\times d}$ of compatible size, the final input matrix, $X_{I}=[x_{{\cal G}};X_{E}]+P_{E}$. The matrix $X_{I}$ is layer normalized and further projected on the row spaces of three learnable weight matrices, $W_{Q},W_{V},W_{k}\in\mathbb{R}^{d\times d}$ to generate the query, key, and value matrices, $Q,K,$ and $V$, respectively, via $Q=X_{I}W_{Q},K=X_{I}W_{K},\;{\rm and}\;V=X_{I}W_{V}.$ These matrices are further divided into $h$ partitions (also, called heads), ${Q^{i},K^{i},V^{i}\in\mathbb{R}^{N\times\tfrac{d}{h}}}$, such that each partition generates a self-attention matrix for the $i^{\rm th}$ head, ${{\cal A}^{i}=\tfrac{Q^{i}{K^{i}}^{\top}}{\sqrt{d/h}}.}$ Finally, we project $V^{i}$ onto the column space of $\sigma({\cal A}^{i})$ as $\sigma({\cal A}^{i})V^{i}$, where $\sigma$ is the Softmax operator applied row-wise. The outputs of different heads are further concatenated into a large matrix ${\begin{bmatrix}\sigma({\cal A}^{1})V^{1}&\sigma({\cal A}^{2})V^{2}&\cdots&\sigma({\cal A}^{h})V^{h}\end{bmatrix}W^{O},}$ post-multiplied by the output weight matrix $W^{O}\in\mathbb{R}^{d\times d}$. This process is sequentially performed in $L$ encoder blocks; see Figure 2(a).
3 How Can We Design Learnable Attention?
In this Section, we design learnable multi-head self-attention (MHSA) module for ViTs that operate inside its encoder blocks.
Last year witnessed a surge in embedding KANs in different DNN architectures (Genet & Inzirillo, 2024; Ferdaus et al., 2024; Li et al., 2024). To our knowledge, the work closely related to ours is KAT, where (Yang & Wang, 2025) replace the MLP layers in the ViTs with KAN layers; also, see (Chen et al., 2024). We later show that deploying learnable attention is a complicated and much-involved process.
Let ${\cal A}^{i,j}\in\mathbb{R}^{N\times N}$ be the attention matrix for $i^{\rm th}$ head in the $j^{\rm th}$ encoder block. For $k\in[N],$ the softmax activation function, $\sigma:\mathbb{R}^{N}\to(0,1)^{N}$, operates on the $k^{\rm th}$ row vector, ${\cal A}^{i,j}_{k,:}\in\mathbb{R}^{1\times N}$ of ${\cal A}^{i,j}$ to produce component-wise output $\sigma({\cal A}^{i,j}_{k,l})=\tfrac{e^{{\cal A}^{i,j}_{k,l}}}{\sum\limits_{n=1}^{N}e^{{\cal A}^{i,j}_{k,n}}}.$ Instead of using the softmax function, we can use a learnable activation function, ${\tilde{\sigma}}$ on the row vectors of each attention head ${\cal A}^{i,j}$. With any choice of the basis functions (e.g., B-Spline, Fourier, Fractals, Wavelets, etc.), the activated attention row vector, ${\tilde{\sigma}(\mathcal{A}^{i,j}_{k,:})}$ for $k\in[N]$ can be written as
$\textstyle{\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]}$ $=\textstyle{\begin{pmatrix}\phi_{11}(\cdot)&\phi_{12}(\cdot)&...&\phi_{1N}(\cdot)\\ \phi_{21}(\cdot)&\phi_{22}(\cdot)&...&\phi_{2N}(\cdot)\\ \vdots&\vdots&\ddots&\vdots\\ \phi_{N1}(\cdot)&\phi_{N2}(\cdot)&...&\phi_{NN}(\cdot)\end{pmatrix}}\scriptscriptstyle{\begin{pmatrix}{\cal A}^{i,j}_{k,1}\\ {\cal A}^{i,j}_{k,2}\\ \vdots\\ {\cal A}^{i,j}_{k,N}\end{pmatrix}}$
$=\textstyle{\left(\Phi^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\right)^{\top}},$ (1)
where $\Phi^{i,j}=[\phi^{i,j}_{pq}]\in\mathbb{R}^{N\times N}$ and each matrix entry, $\phi_{pq}$, is referred to as a learnable unit, represented using a set of basis functions $\{\psi^{i,j}_{m}\}_{m=1}.$ The coefficients associated with the basis functions are called the learnable parameters. Note that, in our convention, ${{\cal A}^{i,j}_{k,:}\in\mathbb{R}^{1\times N}}$ is a row vector. We apply the transpose operation to each row vector, to adopt the convention used in KAN layers. As a result, ${\Phi^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\in\mathbb{R}^{N\times 1},}$ and we transpose it to obtain learnable attention row vectors ${\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]}$; see Figures 2(b)-(c).
Projection on to the probability simplex. The softmax function acts on each row of the attention matrix to create a probability distribution; the learnable attention does not ensure that. To ensure that each row vector of the learned attention matrix lies on a probability simplex, we project them onto the $\ell_{1}$-unit ball. That is, for each attention matrix, ${\cal A}^{i,j}\in\mathbb{R}^{N\times N}$, and for each $k\in[N],$ we want to have ${\|\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]\|_{1}=1}$ and $\tilde{\sigma}\left[({\cal A}^{i,j}_{k,l})\right]\geq 0$ for $l\in[N]$. In the following, we cast it as a sparse approximation problem, whose variants have been well-studied in the past decade and arise frequently in signal processing (Bryan & Leise, 2013; Donoho, 2006; Dutta, 2016) and matrix approximation (Boas et al., 2017; Stewart, 1993).
A sparse approximation problem. Let $y\in\mathbb{R}^{m}$ be a given vector. If we want to approximate $y$ with a vector, $x\in\mathbb{R}^{m}$, with positive components and $\|x\|_{1}=1$, we can write the constrained optimization problem as:
${}x^{\star}=\arg\min_{x\in\mathbb{R}^{m}}\frac{1}{2}\|x-y\|^{2}$
$\text{subject to\;\;}x_{i}\geq 0\;\text{for}\;i\in[m],\text{and}\;\|x\|_{1}=1.$ (2)
For completeness, we describe the solution to problem (3) in §A. Algorithm 1 provides a pseudocode for $\ell_{1}$ projection, and by setting $y=\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]$ and $m=N$ for each $k\in[N]$ in Algorithm 1, we obtain the projected attention vector.
4 Our Architecture
Section 3 outlines a generic learnable attention module. However, we experience computation bottleneck in implementing it. As in (3), for learnable activation function, $\tilde{\sigma}:\mathbb{R}^{N}\to\mathbb{R}^{N}$, an operator matrix, $\Phi^{i,j}\in\mathbb{R}^{N\times N}$ acts on each row ${\cal A}^{i,j}_{k,:}$ of ${\cal A}^{i,j}$, and produces the output $\tilde{\sigma}({\cal A}^{i,j}_{k,:}).$ These operations are extremely compute-heavy and have large memory footprints. E.g., ViT-tiny (Dosovitskiy et al., 2020; Wu et al., 2022; Steiner et al., 2022) has 5.53M parameters and requires nearly 0.9 GB of GPU memory to train. Implementing the learnable attention (3) in ViT-tiny training with B-splines of order 5 and grid size 10 and Fourier basis with grid size 10 increases the parameter count to 30.68M and 39.06M, respectively. We could not train the version with B-splines for its humongous memory requirements, and B-spline computations are non-parallelizable. The one with Fourier basis is parallelizable but it takes approximately 60 GB of GPU memory when computing with a batch of one image. The computational bottleneck is agnostic of the basis functions.
(a) Attention matrix ${\mathcal{A}^{i,j}}$ before softmax activation.
(b) Attention matrix $\sigma(\mathcal{A}^{i,j})$ after softmax activation.
Figure 3: Spectral analysis of the attention matrices before and after softmax shows that they are low-rank. For this experiment, we use all 3 heads in the last encoder block of ViT tiny on 5 randomly sampled images from the CIFAR-10 validation set. We plot all 15 singular vectors (each of 197 dimensions) where the singular values are arranged in non-increasing order.
How can we remedy this? Deep neural networks exhibit low-rank structure (Oja, 1982; Jain et al., 2016). Recently, (Kwon et al., 2024) showed across various networks, the training largely occurs within a low-dimensional subspace. Therefore, we postulate that the attention matrices also have an underlying low-rank structure.
To validate this, we perform spectral analysis of the attention matrices before and after the softmax activation; see Figures 3-3. The scree test (Cattell, 1966) shows that for vanilla attention without softmax, there are 8 significant singular values; after softmax activation this count increases to 16. The above observation verifies that with or without the nonlinear activation, the attention matrix $\mathcal{A}^{i,j}$ possesses an underlying low-rank structure. This motivates us to use a lower-dimensional operator, ${\widehat{\Phi}}$ for the learnable attention calculation with $r=12$. In §B.3, we perform an ablation to find the best $r.$
Instead of using an ${N\times N}$ operator, ${\Phi^{i,j}}$, we use a reduced sized operator ${\widehat{\Phi}^{i,j}\in\mathbb{R}^{r\times N}}$ such that ${r\ll N}$, and the new learned activation is $r$-dimensional for $k\in[N]$. That is, ${\widehat{\Phi}^{i,j}}$ down-projects each attention row vector of ${\cal A}^{i,j}$ to a lower dimensional subspace. This process significantly reduces the computational overhead. Next, we post-multiply using another learnable weight matrix, $W^{i,j}\in\mathbb{R}^{N\times r}$ to project them back to their original dimension. For each $k\in[N]$, this operation results in computing ${\widehat{\sigma}({\cal A}^{i,j}_{k,:})=\left[W^{i,j}\widehat{\Phi}^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\right]^{\top}}$; see Figure 2(d).
Figure 4: Different configurations to update $\widehat{\Phi}$:(a) Blockwise configuration, where $\Phi^{i,1}\neq\Phi^{i,2}\neq\cdots\neq\Phi^{i,L}$ for all $i=1,2,...,h$ (b) universal configuration, where $\Phi^{i,1}=\Phi^{i,2}=\cdots=\Phi^{i,L}=\Phi^{i}$ for all $i=1,2,...,h.$
Fourier Kolmogorov-Arnold Attention. Although the default basis for KANs is B-Splines in (Liu et al., 2025), the number of MHSA parameters is much more than that of an MLP. Specifically, for $L$ encoder blocks, each with $h$ attention heads, the parameter complexity is $O(N^{2}Lh(G+k))$, for $\Phi\in\mathbb{R}^{N\times N}$ and $O(NrLh(G+k))$, for $\widehat{\Phi}\in\mathbb{R}^{r\times N}$, if the model uses B-Splines of degree $k$.
Moreover, as (Yang & Wang, 2025) mentioned, B-splines are localized functions, not standard CUDA functions. Although efficient CUDA implementations for cubic B-Splines exist (Ruijters & Thévenaz, 2012; Ruijters et al., 2008), their overall implementation results in slower, sparse, non-scalable, and complicated GPU execution; also, see (Xu et al., 2024; Pal & Das, 2024). In §B.7, our extensive experiments with B-spline basis in KAN for image classification tasks verify their poor generalizability on medium-scale datasets (e.g., CIFAR-10 and CIFAR-100).
So, what could be an attractive basis? A fundamental question in function approximation is whether the function converges pointwise almost everywhere. Carleson in 1966 proved the following fundamental result for the Fourier approximation of $L^{p}$ periodic functions.
Theorem 2.
(Carleson, 1967) Let $f$ be an ${L^{p}}$ periodic function for ${p\in(1,\infty]}$, with Fourier coefficients ${\hat{f}(n)}$. Then ${\lim_{N\to\infty}\sum_{|n|\leq N}{\hat{f}}(n)e^{inx}=f(x)},$ for almost every $x$.
In particular, if a function is continuously differentiable, its Fourier series converges to it everywhere. In the past, (Xu et al., 2024; Mehrabian et al., 2024) used Fourier basis in KAN, (Dong et al., 2024) proposed Fourier analysis network (FAN) to capture periodic phenomena. Motivated by these, we use the Fourier basis to approximate the effect of the smooth softmax function. For our architecture, we employ the Fourier basis, $\{\sin{(\cdot)},\cos{(\cdot)}\}$ with gridsize $G$ to design learnable attention. Algorithm 1 can be used optionally. For the operator ${\widehat{\Phi}^{i,j}=[\widehat{\phi}^{i,j}_{pq}]\in\mathbb{R}^{r\times N}}$, we have ${\widehat{\Phi}^{i,j}:\mathbb{R}^{N}\to\mathbb{R}^{r}}$. Hence, each row ${{\cal A}^{i,j}_{k,:}}$, with ${k\in[N]}$, transformed into ${\widehat{\Phi}^{i,j}_{p}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]=\sum_{q=1}^{N}\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})},$ via the Fourier bases such that
${\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})=\sum_{m=1}^{G}a_{pqm}\cos{(m{\cal A}^{i,j}_{k,q})}+b_{pqm}\sin{(m{\cal A}^{i,j}_{k,q})},}$
where $m$ refers to the ${m^{\rm th}}$ grid point on a uniform grid of size $G$. The weights, [{${[\{[a_{pqm}],[b_{pqm}]\}_{m=1}^{G},W^{i,j}]}$ are updated in the backpropagation. If we use a Fourier basis in ${\widehat{\Phi}\in\mathbb{R}^{r\times N}}$, the parameter complexity of MHSA becomes ${O(2NrhGL)}.$ Regardless of the basis used, the complexity of linear projection is $O(N^{2}rhL).$ We note that for any Transformer, $L$ and $h$ are fixed. So, one may omit them from the complexity results.
Figure 5: Training loss and test accuracy of vanilla ViTs and their Fourier KArAt versions on CIFAR-10, CIFAR-100 and ImageNet-1K.
Blockwise and Universal operator configuration. We consider two configurations for updating the operator $\widehat{\Phi}^{i,j}$. (a) Blockwise: In this configuration, each encoder block involves learning the attention through $h$ distinct operators $\widehat{\Phi}$ for each of the $h$ heads, totaling $hL$ operators; see Figure 4 (a). Like the MHSA architecture in ViTs, the blockwise configuration is designed to learn as many different data representations as possible. (b) Universal: The motivation behind this configuration comes from the KAT (Yang & Wang, 2025). In KAT, the MLP head is replaced with different variations of a KAN head—KAN and Group-KAN. In Group-KAN, the KAN layer shares rational base functions and their coefficients among the edges. Inspired by this, in our update configuration, all $L$ encoder blocks share the same $h$ operators; $\widehat{\Phi}^{i,j}=\widehat{\Phi}^{i}$ for $j=1,2,\ldots,L$; see Figure 4 (b). Rather than learning attention through $hL$ operators, this configuration only uses $h$ operators. Here, we share all learnable units and their parameters in each head across $L$ blocks. We postulate that blockwise mode with more operators captures more nuances from the data. In contrast, the universal mode is suitable for learning simpler decision boundaries from the data. We also note that these two modes can be used with (3) as shown in Figure  2. Finally, Algorithm 2 in Appendix gives pseudocode of applying Fourier-KArAt in ViT encoder block.
Table 1: Performance of the best-performing Fourier KArAt models compared to the conventional vanilla baselines. The best and the second best Top-1 accuracies are given in red and blue, respectively. The $\downarrow$ and $\uparrow$ arrows indicate the relative loss and gain, respectively, compared to the base models.
Model CIFAR-10 CIFAR-100 Imagenet-1K Parameters
Acc.@1 Acc.@5 Acc.@1 Acc.@5 Acc.@1 Acc.@5
ViT-Base 83.45 99.19 58.07 83.70 72.90 90.56 85.81M
+ $G_{1}B$ 81.81 (1.97$\downarrow$) 99.01 55.92(3.70$\downarrow$) 82.04 68.03(6.68$\downarrow$) 86.41 87.51M (1.98% $\uparrow$)
+ $G_{1}U$ 80.75(3.24$\downarrow$) 98.76 57.36 (1.22% $\downarrow$) 82.89 68.83 (5.58$\downarrow$) 87.69 85.95M (0.16% $\uparrow$)
ViT-Small 81.08 99.02 53.47 82.52 70.50 89.34 22.05M
+ $G_{3}B$ 79.78 (1.60$\downarrow$) 98.70 54.11 (1.20$\uparrow$) 81.02 67.77 (3.87$\downarrow$) 87.51 23.58M (6.94$\uparrow$)
+ $G_{3}U$ 79.52(1.92$\downarrow$) 98.85 53.86(0.73$\uparrow$) 81.45 67.76(3.89$\downarrow$) 87.60 22.18M (0.56$\uparrow$)
ViT-Tiny 72.76 98.14 43.53 75.00 59.15 82.07 5.53M
+ $G_{3}B$ 76.69(5.40$\uparrow$) 98.57 46.29(6.34$\uparrow$) 77.02 59.11 (0.07$\downarrow$) 82.01 6.29M (13.74$\uparrow$)
+ $G_{3}U$ 75.56 (3.85$\uparrow$) 98.48 46.75(7.40$\uparrow$) 76.81 57.97(1.99$\downarrow$) 81.03 5.59M (1.08$\uparrow$)
5 Empirical Analysis
Baselines and Datasets. For our benchmarking, we chose three popular vision Transformers—ViT-tiny, ViT-small (Wu et al., 2022; Steiner et al., 2022), and ViT-base (Dosovitskiy et al., 2020). We incorporate Fourier-KArAt in them by replacing their softmax function with learnable activation. We perform our benchmarking on CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009) and ImageNet-1K (Deng et al., 2009) datasets.
Implementing Fourier KArAt. Our implementation of MHSA using the Fourier basis has one hyperparameter, the grid size, $G$, and two configurations, blockwise or universal. Using different permutations of these factors, we can design multiple variants of Fourier-KArAt. We vary our grid sizes from the set ${G=\{1,3,5,7,9,11\}.}$ For each grid value, we perform universal and blockwise updates; see Figure 4. With this formalization, $G_{n}B$ denotes Fourier-KArAt with grid size $n$ and weights updated in blockwise mode, and $G_{n}U$ denotes Fourier-KArAt with grid size $n$ and weights updated in universal mode. See the detailed ablation study of these different variants in the Appendix. We also note that projecting the learned attention rows to a probability simplex degrades the performance; see §B.5. We dispense $\ell_{1}$ projection for the experiments in the main paper. See implementation details in §B.1.
Figure 6: Weight distribution of (top to bottom) ViT-Tiny, ViT-Tiny+Fourier KArAt, ViT-Base, ViT-Base+Fourier KArAt. The columns (left to right) represent weights at initialization, at epoch 50, at epoch 100, and their superposition.
Figure 7: Spectral decomposition of the attention matrix for ViT-Tiny on CIFAR-10 dataset with vanilla softmax attention and our learnable Fourier attention.
Figure 8: Loss landscape for ViT-Tiny and ViT-Base (the smallest and the largest model) along the two largest principal component directions of the successive change of model parameters. The top row provides a 3D-visulaization of the loss surface including the local optima and saddle points. The bottom row shows the loss contours along with the trajectory of the optimizers.
5.1 Training Performance and Model Quality
Table 1 presents the performance of the best-performing Fourier KArAt variants. We also present the training loss and test accuracy curves on all three datasets; see Figure 5. Fourier KArAt is easily optimized in the initial training phase, gaining at par or better loss value and accuracy than the vanilla ViTs. However, in the later training phase, except for the fewer parameter variants (ViT Tiny+Fourier KArAt), the loss curve flattens faster. From Figure 5 we infer that for the majority of the scenarios involving Fourier-KArAt. We also notice that larger models like ViT-Base are more susceptible to slight changes in the parameters, making the training process difficult to manage, which we try to investigate in the rest of the paper using several analytical studies. To conclude, ViT-Tiny+Fourier KArAt variants outperform ViT-Tiny on CIFAR-10 and CIFAR-100 by $5.40\%$ and $7.40\%$, respectively, and on ImageNet-1K it achieves almost similar accuracy. ViT-Small and -Base models with Fourier KArAt variants can barely outperform the vanilla models and the accuracy differences increase.
Comparing the Top-1 accuracies from Table 1 across blockwise and universal modes on the datasets does not have enough evidence to support the claim that one mode is better than the other. These tabulated accuracies are the results of applying the two variants on only two grid sizes.
Figure 9: Visualizing attention maps for Vit-Tiny. The original images used for inference are on the left, and on the right, we show the attention score map and the image regions of the dominant head (Top row: Fourier KArAt, bottom row: Vanilla MHSA).
5.2 Analysis of Performance
In this Section, we dissect the performance of Fourier KArAt through a diverse set of analyses.
(i) Distribution of the Weights. (Zhang et al., 2022) considered an invariant measure perspective to describe the training loss stabilization of the neural networks. Based on this idea, we study the distribution of the weights of the smallest and largest models of the ViTs, Tiny and Base, and their Fourier KArAt variants. Figure 6 shows the distribution of the weights of these models during different training phases. The evolution of the weights’ distributions for respective models remains invariant, and we can guarantee the convergence of their loss values; also, see Figure 5. However, as explained by (Zhang et al., 2022), this perspective cannot shed light on the generalization capacity and structural differences between different neural networks.
(ii) Spectral Analysis of Attention. Although by studying the distribution of the weights we cannot comment on the generalizability, from Table 1 we realize all the KArAt variants have more parameters than their vanilla counterparts. Therefore, it would be interesting to see how their attention matrices behave. As discussed in §4, the attention matrices in vanilla MHSA have a low-rank structure. Following that study, we investigated if Fourier KArAt’s learned attention matrices are also low-rank. We use all 3 heads in the last encoder block of ViT-Tiny on 5 randomly sampled images from the CIFAR-10 validation set. There are a total of 15 singular vectors (each of 197 dimensions) where the singular values are arranged in non-increasing order. Let $\sigma_{i}$ be the $i^{\rm th}$ singular value across all heads and samples (it is permutation invariant). For each $i\in[197],$ we plot $[\ln(\sigma_{i}^{\min}),\ln(\bar{\sigma_{i}}),\ln(\sigma_{i}^{\max})]$, where $\bar{\sigma_{i}}$ is the average of $i^{\rm th}$-indexed singular value across all samples and heads. From Figure 7, we observe that vanilla softmax attention has significantly larger singular values than its KArAt variant, $G_{3}B$. Fourier KArAt with more parameters shows a faster drop in significant singular values. This observation indicates that Fourier KArAt variants are highly overparameterized and have a much lower rank structure compared to the vanilla ViTs.
(iii) Loss Landscape Analysis. To understand the reason behind the slower convergence in the later training phase and the lower generalizability of KArAts, we visualize the loss landscape of the networks. We plot the loss surface along the directions in which the gradients converge. Following (Li et al., 2018), we perform principal component analysis (PCA) of the change in parameters over training progression to understand the major directions of parameter convergence. Considering the fully trained model as the minimum in the loss hyperplane and the two principal component directions as $X$ and $Y$ axes, we plot the loss values over the validation set of CIFAR-10 along the $Z$ axis for ViT-Tiny and ViT-Base for the vanilla MHSA and KArAt ($G_{3}B$ and $G_{1}B$, respectively). These visualizations in Figure 8 show that having Fourier KArAt in ViT architectures significantly impacts the smoothness of the loss surfaces.
ViT-Tiny with the least parameters has the smoothest loss landscape and modest generalizability. In contrast, ViT-Tiny+Fourier KArAt’s loss landscape is spiky; it indicates the model is full of small-volume minima (Huang et al., 2020). However, the model is modest in the number of parameters, so the gradient descent optimizer can still find an optimized model with better generalizability than the vanilla ViT-Tiny; hence it gains a better test accuracy; see Table 1. ViT-Base, however, has more parameters than Tiny and its loss surface is much spikier than ViT-Tiny. Finally, the loss surface of ViT-Base+Fourier KArAt is most spikey, making it a narrow margin model with sharp minima in which small perturbations in the parameter space lead to high misclassification due to their exponentially larger volume in high dimensional spaces; see Figure 8(d). Moreover, ViT-Base+$G_{3}B$ has 14 times more parameters than ViT-Tiny+$G_{1}B$. With learnable activations, gradient descent optimizers fail to find the best-optimized model, as small differences in the local minima translate to exponentially large disparities. The slow convergence in the later phases of the models is bacause the increasing number of local minima can slow down convergence.
(iv) Attention Visualization. MHSA in ViTs captures region-to-region interaction in images. DINO (Caron et al., 2021) provides an innovative way to explain the dominant regions in an image that contribute towards the inference decision. It maps the self-attention values of the CLS token at the last encoder layer which is directly responsible for capturing the information related to the class label. While the computer vision community predominantly uses this attention visualization, it is not directly suitable for our case. Unlike softmax, Fourier KArAt does not inherently ensure learned attention values in $[0,1]$, as we dispense the $\ell_{1}$ projection. Therefore, we cannot interpret its performance similarly. Nonetheless, pre-softmax or pre-KArAt values in attention matrix $\mathcal{A}^{i,j}$ are supposed to capture token-to-token interactions, and we adapt the attention maps to ignore the negative values for the sake of visualization. We use the ImageNet-1K trained models of ViT-Tiny (with vanilla MHSA and Fourier KArAt) and plot the attention maps of the last layers in Figure 9. Fourier KArAts tend to concentrate high interaction scores on the concerned object’s dominant features compared to having interactions throughout the object regions like traditional ViTs (Dosovitskiy et al., 2020; Touvron et al., 2021).
6 Conclusion and Future Direction
This paper proposes a learnable MHSA for vision Transformers that can theoretically operate on any choice of basis. The central design theme of this learnable attention module is based on the famous Kolmogorov-Arnold representation theorem (Kolmogorov, 1956) and the Kolmogorov-Arnold network proposed by (Liu et al., 2025).
This endeavor raises many interesting questions regarding KArAt’s computing scalability and the resulting models’ generalization capacity. In the future, with the advent of large multimodal models (LMMs), we plan to extend this study beyond ViTs for image classification and check the resilience of KArAt in the language processing domain and on open-source LMMs.
KANs are parameter efficient—this claim does not hold for learnable KArAt selft-attention design. However, our spectral analysis and loss-landscape experiments suggest that overparameterized KArAts’ weights lie in a much lower-dimensional subspace. Nevertheless, how to navigate through KArAts’ parameter-heavy loss landscape to find the optimized model with better generalizability remains an open question that we encourage the community to focus on. Finally, we note that the data specificity of MHSA plays a key role in their high performance. Therefore, we aim to explore its scope with KArAts with the possibility of pitfalls due to long-range dependency (Park & Kim, 2022).
Appendix A Solution to Problem (3)
First, we rewrite the constrained problem (3) as an unconstrained problem using the Lagrange multipliers as:
${\cal L}(x,\lambda,\mu)=\frac{1}{2}\|x-y\|^{2}+\lambda\bigg{(}\sum\limits_{i=1}^{m}x_{i}-1\bigg{)}-\mu^{T}x,$ (3)
where $\lambda,\mu$ are Lagrange multipliers. Using the Karush-Kuhn-Tucker (KKT) stationarity condition on (3), we find $0\in\partial{\cal L}(x,\lambda,\mu)$, which implies, $x_{i}-y_{i}+\lambda=\mu_{i}.$ The complementary slackness gives, $\mu_{i}x_{i}=0$ for $i\in[m].$ Further, for $i\in[m]$, the primal and dual feasibility conditions are $x_{i}\geq 0,$ $\sum\limits_{i=1}^{m}x_{i}=1$, and $\mu_{i}\geq 0,$ respectively.
The stationarity and complementary slackness conditions give $x_{i}-y_{i}+\lambda\geq 0,\ (x_{i}-y_{i}+\lambda)x_{i}=0$. If $\lambda\geq y_{i}$, then $x_{i}=0$. Otherwise, we have that $\sum\limits_{i=1}^{m}x_{i}=\sum\limits_{i=1}^{m}(y_{i}-\lambda)=1$ implying $\lambda=\frac{1}{m}(\sum\limits_{i=1}^{m}y_{i}-1)$.
Algorithm 2 Fourier-KArAt in $j^{\rm th}$ Vision Transformer block
  Input: $X\in\mathbb{R}^{N\times d},\text{ Fourier basis }\{\sin{(\cdot)},\cos{(\cdot)}\}$
  Output: $O\in\mathbb{R}^{N\times d}$
  Parameters: $W_{Q}^{i},W_{K}^{i},W_{V}^{i}\in\mathbb{R}^{d\times d_{h}},d_{h}=\frac{d}{h},G,[\{[a_{pqm}],[b_{pqm}]\}_{m=1}^{G},W^{i,j}],W^{O,j}$
  Hyperparameters: Blockwise, universal, Algorithm 1 ($\ell_{1}$ projection), $L$ encoder blocks
  for each head $i\in\{1,\ldots,h\}$ do:
         $Q^{i,j}\leftarrow XW^{i,j}_{Q}$
         $K^{i,j}\leftarrow XW^{i,j}_{K}$
         $V^{i,j}\leftarrow XW^{i,j}_{V}$
         ${\cal A}^{i,j}=\frac{Q^{i,j}(K^{i,j})^{\top}}{\sqrt{d_{h}}}$
  end for
  for each head $i\in\{1,\ldots,h\}$ do:
         for each row $k\in\{1,\ldots,N\}$ do:
              $\widehat{\Phi}^{i,j}[({\cal A}^{i,j})_{k,:}^{\top}]\leftarrow\sum_{q=1}^{N}\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})=\sum_{m=1}^{G}a_{pqm}\cos{(m{\cal A}^{i,j}_{k,q})}+b_{pqm}\sin{(m{\cal A}^{i,j}_{k,q})}$
              if  $\ell_{1}$ projection then
                   Execute Algorithm 1
              else
                   pass
              $\widehat{\sigma}({\cal A}^{i,j}_{k,:})\leftarrow\left[W^{i,j}\widehat{\Phi}^{i,j}[({\cal A}^{i,j}_{k,:})^{\top}]\right]^{\top}$
         end for
  end for
  $O\leftarrow\begin{bmatrix}\widehat{\sigma}({\cal A}^{1,j})V^{1,j}&\widehat{\sigma}({\cal A}^{2,j})V^{2,j}&\cdots&\widehat{\sigma}({\cal A}^{N,j})V^{N,j}\end{bmatrix}W^{O,j}$
  if Blockwise Mode then
         Return $O$
  else if universal Mode then
         pass learnable units and parameters to $(j+1)^{\rm th}$ encoder block
         Return $O$
Note that, in Algorithm 1 one can replace $y$ with $\widehat{\Phi}^{i,j}[({\cal A}^{i,j})_{k,:}^{\top}]$ so that the constraint ${\|\widehat{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]\|_{1}=1}$ and $\widehat{\sigma}\left[({\cal A}^{i,j}_{k,l})\right]\geq 0$ is satisfied for the low-rank attention matrices in Fourier-KArAt. This algorithm is also kept as an optional sub-process for Algorithm 2.
Appendix B Addendum to Empirical Analysis
This section complements our empirical results in §5.
B.1 Implementation Details
We implement the framework in Python using PyTorch (Paszke et al., 2019) and use the training strategy of DEiT (Touvron et al., 2021). We train all models for $100$ epochs with the ADAM optimizer (Kingma & Ba, 2015) with a base learning rate of $3.125\times 10^{-5}$ and batch-size $32$, except for the experiments on ImageNet-1K that use a learning rate of $1.25\times 10^{-4}$ with batch-size $128$. All experiments use a warm-up of $5$ epochs and a cosine scheduler with a weight decay of $5\times 10^{-2}$. The experiments were performed on two 80 GB NVIDIA H100 GPUs. All the hyper-parameter settings are given in Table 2.
Table 2: Hyper-parameter settings for all the experiments conducted in this work.
Input Size $224\times 224$
Crop Ratio $0.9$
Batch Size 128 for Imagenet-1K and 32 for CIFAR-10 & CIFAR-100
Optimizer AdamW
Optimizer Epsilon $1\times 10^{-6}$
Momentum $0.9$
Weight Decay $0.05$
Gradient Clip $1.0$
Learning Rate Schedule Cosine
Learning Rate $5\times 10^{-4}\times\frac{\text{Batch Size}}{512}$
Warmup LR $1\times 10^{-6}$
Min LR $1\times 10^{-5}$
Epochs 100
Decay Epochs 1
Warmup Epochs 5
Decay Rate $0.988$
Exponential Moving Average (EMA) True
EMA Decay $0.99992$
Random Resize & Crop Scale and Ratio $(0.08,1.0),(0.67,1.5)$
Random Flip Horizontal $0.5$; Vertical $0.0$
Color Jittering $0.4$
Auto-augmentation rand-m15-n2-mstd1.0-inc1
Mixup True
Cutmix True
Mixup, Cutmix Probability $0.5$, $0.5$
Mixup Mode Batch
Label Smoothing $0.1$
B.2 Computation Time and Throughput
The overall computation for Fourier KArAt variants is higher than their conventional softmax MHSA, and we have delineated the same in Figure 10. Primarily, the Fourier KArAt variants have a longer training time. We show the training time comparison between the vanilla MHSA and its Fourier KArAt versions for 100 epochs on CIFAR-10, CIFAR-100, and Imagenet-1K datasets for all the models (ViT-Tiny, ViT-Small, ViT-Base) in Figures 101010, respectively. We only compare the best-performing Fourier KArAt models ($G_{1}B$ and $G_{1}U$ for ViT-Base, and $G_{3}B$ and $G_{3}U$ for ViT-Tiny & ViT-Small) with their softmax MHSA counterparts. We also observe that universal mode $G_{n}U$ training times are consistently slightly less than the blockwise modes $G_{n}B$.
During the training, we monitored the GPU memory requirements, and as expected, Fourier KArAt variants utilize significantly more memory than vanilla MHSA. In particular, the GPU memory requirements scale by 2.5×2.5$\times$2.5 ×, 4.25×4.25$\times$4.25 ×, and 6×6$\times$6 × in case of ViT-Tiny, ViT-Small, and ViT-Base, respectively, compared to the vanilla softmax MHSA.
We also compare the throughput during inference in Figure 10 and see slightly faster inference in universal mode than blockwise except for ViT-Base. While there is a massive training time discrepancy between vanilla ViTs and Fourier KArAt ViTs, the inference speeds for Fourier KArAt variants are comparable to their vanilla counterparts. Although there is a minor difference in throughput between universal mode and blockwise mode during inference, theoretically, both variants for any model with the same grid size should involve the same quantity of FLOPs.
(a) Training time on CIFAR-10 dataset.
(b) Training time on CIFAR-100 dataset.
(c) Training time on Imagenet-1K dataset.
(d) Throughput comparison during inference.
Figure 10: A detailed comparison of computing requirements. We compare the training times for 100 epochs with the hyperparameter settings given in Table 2 for all the datasets CIFAR10, CIFAR100, and ImageNet1K. We also compare the throughputs of different models on ImageNet1k; the throughput results will be similar for other datasets as the input size is $224\times 224$.
B.3 Ablation with the Hidden Dimension, $r$
While avoiding the computational overhead for computing $\Phi^{i,j}\in\mathbb{R}^{N\times N}$, we make use of the low-rank structure that the attention heads show (see Figure 3) by comparing different values of $r$. Particularly, we consider the values, $r=8,12,24$, on the Fourier KArAt variant of ViT-Base model, $\widehat{\Phi}^{i,j}\in\mathbb{R}^{r\times N}$; see Table 5 for results on CIFAR-10. In this ablation, we observe that changing the hidden dimension has a minimal impact on the model’s performance. This can be explained by the sudden drop in the singular values, as shown in Figure 7. As long as $r$ remains greater than the sudden drop index, the model should not be impacted by the changing of $r$ except for changes in computational requirement; a higher $r$ would incur a higher compute time as the size of the operator $\widehat{\Phi}$ scales with $r$.
Table 3: Ablation on grid size $G$ for ViT-Tiny, ViT-Small and ViT-Base. We find a particular grid size suitable for each of the models.
     Model      Acc.@1
     CIFAR-10      CIFAR-100
     ViT-Base      83.45      58.07
     + $G_{1}B$      81.81      55.92
     + $G_{1}U$      80.75      57.36
     + $G_{3}B$      80.09      56.01
     + $G_{3}U$      81.00      57.15
     + $G_{5}B$      79.80      54.83
     + $G_{5}U$      81.17      56.38
     + $G_{11}B$      50.47      42.02
     + $G_{11}U$      40.74      39.85
     ViT-Small      81.08      53.47
     + $G_{1}B$      79.00      53.07
     + $G_{1}U$      66.18      53.86
     + $G_{3}B$      79.78      54.11
     + $G_{3}U$      79.52      53.86
     + $G_{5}B$      78.64      53.42
     + $G_{5}U$      78.75      54.21
     + $G_{11}B$      77.39      52.62
     + $G_{11}U$      78.57      53.35
     ViT-Tiny      72.76      43.53
     + $G_{1}B$      75.75      45.77
     + $G_{1}U$      74.94      46.00
     + $G_{3}B$      76.69      46.29
     + $G_{3}U$      75.56      46.75
     + $G_{5}B$      75.85      –
     + $G_{5}U$      74.71      –
     + $G_{7}B$      75.11      –
     + $G_{7}U$      74.45      –
     + $G_{9}B$      74.85      –
     + $G_{9}U$      73.97      –
     + $G_{11}B$      74.52      –
     + $G_{11}U$      73.58      –
Table 4: Ablation on hidden dimension $r$ on CIFAR-10 with ViT-Base. Here, we compare the values of $r\in\{8,12,24\}$.
Model $r$ Acc.@1 on CIFAR-10
ViT-Base 58.07
+ $G_{3}B$ 24 80.54
+ $G_{3}U$ 24 80.81
+ $G_{5}B$ 24 77.99
+ $G_{5}U$ 24 80.52
+ $G_{3}B$ 12 80.09
+ $G_{3}U$ 12 81.00
+ $G_{5}B$ 12 79.80
+ $G_{5}U$ 12 81.17
+ $G_{3}B$ 8 80.76
+ $G_{3}U$ 8 80.40
+ $G_{5}B$ 8 79.79
+ $G_{5}U$ 8 80.83
Table 5: Comparing performance of Fourier KArAt for ViT-Tiny and ViT-Base with or without using $\ell_{1}$ projection in Algorithm 1. ViT-Tiny and ViT-Base use softmax and do not require $\ell_{1}$ projection.
Model $\ell_{1}$ Projection Acc.@1
on CIFAR-10
ViT- Tiny 72.76
+ $G_{3}B$ 41.99
+ $G_{3}U$ 40.85
+ $G_{3}B$ 76.69
+ $G_{3}U$ 75.56
ViT-Base 83.45
+ $G_{3}B$ 47.44
+ $G_{3}U$ 46.11
+ $G_{3}B$ 80.09
+ $G_{3}U$ 81.00
B.4 Ablation Study on the Impact of grid size, $G$
KANs are highly dependent on certain hyperparameters, and Fourier KAN and Fourier-KArAt have only one hyperparameter to tune the performance — grid size $G$. Thus, we perform extensive experimentation involving grid size $G$ and present in Table 5. We observe that each of the particular ViT models, in conjunction with particular Fourier-KArAt variants, has a typical $G$ value that brings out its best performance, and there is no universal value of $G$ to follow. When performing validation with ViT-Base+ variants on CIFAR-10 and CIFAR-100, the accuracy drops as the grid size passes a size after 5. However, this behavior is not persistent with the ViT-Small/Tiny+ variants; see Table 5111We did not perform gridsize 5, 7, 9, and 11 experiments with ViT-Tiny+KArAt as it was already outperforming the base ViT-Tiny with gridsize 1 and 3, and the experiments are extensively resource intensive..
Table 6: Ablation study on the operator variants in Fourier KArAt. Note that $\widehat{\Phi}(W{\cal A})$, $W\widehat{\Phi}({\cal A})$ refer to order of using $W$ and $\widehat{\Phi}$, and $\Phi$ refers to the full-rank operator.
Model Acc.@1 on CIFAR-10
$\widehat{\Phi}(W{\cal A})$ $W\widehat{\Phi}({\cal A})$ $\Phi({\cal A})$
ViT-Base
+ $G_{11}B$ 41.02 50.47 30.14
+ $G_{11}U$ 47.74
ViT-Tiny
+ $G_{11}B$ 67.04 75.41 33.09
+ $G_{11}U$ 70.98 73.67 36.98
Table 7: Performance using softmax and variants of Fourier KArAt in separate heads.
Model Low Rank Configurations No. of Distinct Heads Acc.@1 on
(for Fourier KArAt) softmax Fourier KArAt CIFAR-10
ViT-Base 12 0 83.45
ViT-Base + $G_{11}B$ $\hat{\sigma}=\widehat{\Phi}(W{\cal A})$ 0 12 41.02
$\hat{\sigma}=\widehat{\Phi}(W{\cal A})$ 6 6 47.60
$\hat{\sigma}=W\widehat{\Phi}({\cal A})$ 0 12 50.47
$\hat{\sigma}={\Phi}({\cal A})$ 0 12 30.14
$\hat{\sigma}={\Phi}({\cal A})$ 6 6 30.24
ViT-Tiny 3 0 72.76
ViT-Tiny + $G_{11}B$ $\hat{\sigma}={\Phi}({\cal A})$ 0 3 33.09
$\hat{\sigma}={\Phi}({\cal A})$ 1 2 32.51
$\hat{\sigma}={\Phi}({\cal A})$ 2 1 34.73
ViT-Tiny + $G_{11}U$ $\hat{\sigma}={\Phi}({\cal A})$ 0 3 36.98
$\hat{\sigma}={\Phi}({\cal A})$ 1 2 33.12
$\hat{\sigma}={\Phi}({\cal A})$ 2 1 70.57
Table 8: Experimental results on multi-layer KANs organized similarly to an MLP. These experiments involve B-Splines as the basis functions, as mentioned in Section 2.
(a) Experiments on CIFAR-10 and CIFAR-100.
Spline Order Grids Grid Range Layers Acc.@1
CIFAR-10 CIFAR-100
1 10 (-4, 4) (10, 10) 22 1
20 (-4, 4) (10, 10) 21 4
(-6, 6) (10, 10) 22 4
(-8, 8) (10, 10) 22 4
40 (-4, 4) (10, 10) 22 1
(-6, 6) (10, 10) 22 1
(-8, 8) (10, 10) 22 1
3 10 (-4, 4) (10, 10) 22 1
20 (-4, 4) (10, 10) 22 4
(-6, 6) (10, 10) 22 4
(-8, 8) (10, 10) 39 4
40 (-4, 4) (10, 10) 22 1
(-6, 6) (10, 10) 29 1
(-8, 8) (10, 10) 31 1
4 10 (-4, 4) (10, 10) 21 1
5 10 (-1, 1) (10, 10) 10 4
(-2, 2) (10, 10) 10 1
(-4, 4) (10, 10) 22 4
(-4, 4) (20, 20) 40 4
(-4, 4) (40, 40) 45 9
(-6, 6) (10, 10) 39 7
(-8, 8) (10, 10) 42 9
20 (-4, 4) (10, 10) 10 1
40 (-4, 4) (10, 10) 21 1
(b) Experiments on MNIST and Fashion-MNIST.
Spline Order Grids Grid Range Layers Acc.@1
MNIST FMNIST
1 10 (-4, 4) (10, 10) 92 85
3 10 (-4, 4) (10, 10) 94 86
4 10 (-4, 4) (10, 10) 94 86
5 10 (-1, 1) (10, 10) 89 83
(-2, 2) (10, 10) 90 95
(-4, 4) (10, 10) 95 86
(-4, 4) (20, 40) 96 87
(-4, 4) (40, 40) 96 88
(-6, 6) (10, 10) 95 86
(-8, 8) (10, 10) 95 86
20 (-4, 4) (10, 10) 93 85
40 (-4, 4) (10, 10) 91 83
B.5 Does Fourier KArAt Require $\ell_{1}$-Projection?
To ensure that each row vector of the learned attention matrix lies on a probability simplex, we project them onto the $\ell_{1}$-unit ball. We use Algorithm 1 to project learned attention vectors to the probability simplex and compare Top-1 accuracies to the baseline model. We note that using $\ell_{1}$ projection does not increase the training time substantially. However, from Table 5, we observe that incorporating this algorithm in the Fourier-KArAt does not improve its performance; instead, the performance degrades significantly. Without applying Algorithm 1, the weights in each learnable unit can easily adjust and update accordingly.
B.6 Alternate Variants of Fourier KArAt
In this section, we experiment with different attention variants on Fourier KArAt.
B.6.1 Alternative Approaches to the Lower Rank Attention Structures
In the main paper, we approximated the effect of the operator, ${\Phi^{i,j}}\in\mathbb{R}^{N\times N}$ casting it as the product of two rank-$r$ operators—one operator with learnable activation, $\widehat{\Phi},$ and the other is the learnable the linear transformation matrix, $W$ such that $\Phi=\widehat{\Phi}W$. Here, we abuse the notations for simplicity. However, a natural question is how many different configurations are possible with $\widehat{\Phi}$ and $W$ such that they can approximate the effect of $\Phi.$ Specifically, we use the following configurations: (i) [^2] This configuration was used in the main paper. $W\widehat{\Phi}(\mathcal{A}),$ where $\widehat{\Phi}\in\mathbb{R}^{r\times N}$ and $W\in\mathbb{R}^{N\times r}$, and (ii) $\widehat{\Phi}(W{\cal A}),$ where $W\in\mathbb{R}^{r\times N}$ and $\widehat{\Phi}\in\mathbb{R}^{N\times r}$. In the first configuration, $\widehat{\Phi}(\cdot)$ acts on each row of ${\cal A}$ to produce a $r$-dimensional vector, and then $W$ projects it back to an $N$-dimensional subspace. In the second configuration, $W$, first projects each row of ${\cal A}$ to produce a $r$-dimensional vector, and then $\widehat{\Phi}$ with learnable activation produces $N$-dimensional vectors.
Primarily, we started with the full-rank operator $\Phi\in\mathbb{R}^{N\times N}$ and found their calculations are prohibitively expensive, regardless of the choice of the basis. Next, we conduct an ablation study to see which operator configuration works better; see Table 7. Our experiments show that the operator configuration (ii) demonstrates an inferior performance. We postulate that by down-projecting the $N$ dimensional attention row vector to a smaller dimension, $r$ loses adequate token-to-token interaction, and this fails to capture dependencies from significant attention units within a row. After that, from this limited information, learnable activation cannot significantly help the model’s performance.
B.6.2 Fourier KArAt and softmax—An Hybrid version
We consider another variant where we mix the learnable activation and pre-defined activation in each head. E.g., the user can have 2 of the 3 attention heads in each encoder block in ViT-Tiny activated by Softmax and the third activated by the Fourier KArAt. With the idea of incorporating KAN as a way to replace softmax activation, we are curious to see if the ViT model performs better with a hybrid mode of activation; see Table 7 for results.
B.6.3 More Combinations
We have also experimented with more configurations of Fourier KArAt in various permutations of the strategies mentioned above and have not found any significant combination in terms of performance. We also carefully design a particular training strategy where the attention operators ($\widehat{\Phi}$, $W$, and ${\Phi}$ wherever applicable) of Fourier KArAt are trained with a separate learning rate to alleviate the problem of the mismatch of the data passing through these learnable layers, as they consider each row of an attention matrix ${\cal A}$ as an input. This also attends to our observation of the models having gradient and loss explosion during training. However, we found this strategy to be of no improvement over the performance without the same.
B.7 B-Spline experiments
We test the performance of the B-spline KANs in the classification task on the CIFAR10 (Krizhevsky et al., 2009), CIFAR100 (Krizhevsky et al., 2009), MNIST (LeCun et al., 1998, 2010), and Fashion-MNIST (Xiao et al., 2017) datasets. We performed extensive experiments to see if there are any benefits to choosing B-Splines for the basis functions in KAN layers. Table 8 shows the Top-1 accuracies from different variants of a Deep KAN. In this set of experiments, we closely followed earlier works (Yu et al., 2024; Liu et al., 2025). Hyperparameters involved in these experiments include the order of the B-spline $k$, grid size $G$, grid range $[-I,I]$, layers, and width. Although the B-spline KANs yield high accuracies in the small-scale MNIST and Fashion-MNIST datasets, they fail to generalize over larger datasets.