### v3

# Kolmogorov-Arnold Attention: Is Learnable Attention Better For Vision Transformers?

Abstract
Kolmogorov-Arnold networks (KANs) are a remarkable innovation that consists of learnable activation functions, with the potential to capture more complex relationships from data. Presently, KANs are deployed by replacing multilayer perceptrons (MLPs) in deep networks, including advanced architectures such as vision Transformers (ViTs). Given the success of replacing MLP with KAN, this work asks whether KAN could learn token interactions. In this paper, we design the first learnable attention called Kolmogorov-Arnold Attention (KArAt) for ViTs that can operate on any basis, ranging from Fourier, Wavelets, Splines, to Rational Functions. However, learnable activations in the attention cause a memory explosion. To remedy this, we propose a modular version of KArAt that uses a low-rank approximation. By adopting the Fourier basis into this, Fourier-KArAt and its variants, in some cases, outperform their traditional softmax counterparts, or show comparable performance on CIFAR-10, CIFAR-100, and ImageNet-1K datasets. We also deploy Fourier KArAt to ConViT and Swin-Transformer, and use it in detection and segmentation with ViT-Det. We dissect the performance of these architectures on the classification task by analyzing their loss landscapes, weight distributions, optimizer paths, attention visualizations, and transferability to other datasets, and contrast them with vanilla ViTs. KArAt’s learnable activation yields a better attention score across all ViTs, indicating improved token-to-token interactions and contributing to enhanced inference. Still, its generalizability does not scale with larger ViTs. However, many factors, including the present computing interface, affect the relative performance of parameter- and memory-heavy KArAts. We note that the goal of this paper is not to produce efficient attention or challenge the traditional activations; by designing KArAt, we are the first to show that attention can be learned and encourage researchers to explore KArAt in conjunction with more advanced architectures that require a careful understanding of learnable activations. Our open-source code and implementation details are available on: subhajitmaity.me/KArAt
Keywords Multi-Head Self-Attention $\cdot$ Vision Transformers $\cdot$ Kolmogorov-Arnold Network $\cdot$ Kolmogorov-Arnold Transformers $\cdot$ Kolmogorov-Arnold Attention
1 Introduction
Figure 1: (a) Model parameters vs. Top-1 accuracy in ImegeNet-1K training of vanilla ViTs dosovitskiy2020vit , Vision KAN (DeiT+KAN) by VisionKAN2024 , ViT+KAN and Kolmogorov-Arnold Transformer (KAT) by yang2024kolmogorov .(b-i) The traditional softmax attention. (b-ii) The Kolmogorov-Arnold Attention (KArAt) replaces the softmax with a learnable operator, $\Phi^{i,j}$. (b-iii) Regular KArAt uses an operator matrix, $\Phi^{i,j}$ with $N^{2}$ learnable units acting on each row of ${\cal A}^{i,j}$, and is prohibitively expensive. (b-iv) Modular KArAt uses an operator $\widehat{\Phi}^{i,j}\in\mathbb{R}^{N\times r}$ with $r\ll N$, followed by a learnable linear projector $W\in\mathbb{R}^{r\times N}$.
Artificial general intelligence has become a rapidly growing research direction, and Kolmogorov-Arnold Network (KAN) liu2024kan2 marks a remarkable innovation in that. KANs with learnable activation functions can potentially capture more complex relationships and facilitate meaningful interaction between the model and human intuition. After training a KAN on a specific problem, researchers can extract the learned univariate functions that the model uses to approximate complex multivariable functions. By studying these learned functions, researchers can gain insights into the underlying relationships from the data and refine the model. KANs exhibit state-of-the-art performance in finding symbolic function representations yu2024kanmlpfairercomparison , continual learning of one-dimensional functions liu2024kan .
KANs were integrated with neural network architectures, the primary ones being conventional MLPs or convolution neural networks (CNNs). E.g., ferdaus2024kanice ; bodner2024convolutional ; drokin2024kolmogorov ; abueidda2025deepokan ; wang2025intrusion combine KAN with CNNs, li2024UKAN combine KAN with U-Net, etc. Interestingly, (yu2024kanmlpfairercomparison, ) claimed to make the first fairer comparison between KANs and MLPs on multiple ML tasks on small-scale datasets. By setting the same parameter counts in both networks, yu2024kanmlpfairercomparison concluded that KANs with B-Spline basis significantly underperform compared to their MLP counterparts in all ML tasks except symbolic formula representation; azam2024suitability reinstates a similar claim for vision. However, KAN’s exploration of more advanced architectures, such as Transformers, remains limited. VisionKAN2024 replace the MLP layers inside the encoder blocks of data-efficient image Transformers (DeiTs) deit with KANs and proposed DeiT+KAN, yang2024kolmogorov proposed two variants: ViT+KAN that replaces the MLP layers inside ViT’s encoder blocks, and Kolmogorov-Arnold Transformer (KAT), albeit similar to ViT+KAN but with a refined group-KAN strategy.
While KANs were claimed to be not superior to simple MLPs in vision tasks yu2024kanmlpfairercomparison ; azam2024suitability , we were curious to find out how ViTs can see images through the lens of KANs. In that effort, Figure 1(a) demonstrates the performance of vanilla ViTs (-Tiny, -Small wu2022tinyvit ; steiner2022howtotrainvit , and -Base dosovitskiy2020vit ) and their KAN counterparts on ImageNet-1K deng2009imagenet training. While different variants of DeiT+KAN VisionKAN2024 and ViT+KAN yang2024kolmogorov show a 5%–18% drop in the Top-1 accuracy compared to vanilla ViTs and DeiTs in ImageNet-1K, the KATs yang2024kolmogorov show approximately 1%–2.5% gain in Top-1 accuracy compared to vanilla ViTs and DeiTs by keeping about the same number of parameters. These results show KANs require further investigation in more advanced architectures such as Transformers. Nevertheless, first, we want to understand why there is a discrepancy in these two seemingly equivalent implementations.
Searching for an answer, we realized that while DeiT+KAN and ViT+KAN replace the MLP layer with KAN in the Transformer’s encoder block, KAT implements a sophisticated group-KAN strategy that reuses a learnable function among a group of units in the same layer and chooses different bases for different encoder blocks. This strategy improves accuracy by keeping almost the same parameter counts. Therefore, simply replacing MLPs with KANs might not guarantee better performance, but a properly designed KAN could. But, can we only learn attention?
Irrespective of the tasks or datasets at hand, traditional multihead attention, the heart of the Transformers, employs a softmax non-linearity as a one-size-fits-all mechanism for token interactions without the necessary justification for its use. We hypothesize that adopting learnable KAN operators could be a better fit for approximating the underlying non-linear relationship between image patches, providing the flexibility for modeling token interactions in ViTs. Therefore, in the rise of next-generation Transformers, such as Show-o xie2025showo , RL-VLM-F wang2024rlvlmf , Google’s TITAN behrouz2024titans , and SAKANA AI’s Transformer2 transformer2 , that mimic the human brain, we ask: How can we deploy learnable multi-head self-attention to the (vision) Transformers?
We note that there is a line of research that proposes efficient attention mechanisms in terms of sparse attention fibottention ; yun2020n ; shi2021sparsebert ; kovaleva2019BERT ; zhang2021multi ; zaheer2020big ; guo2019star or linear/kernelized attention hanbridging ; nguyen2023probabilistic ; nguyen2021fmmformer ; nguyen2022fourierformer ; lu2021soft ; nguyen2022improving-icml , to remedy computational and memory complexities of attention calculation. In contrast, this paper tries to understand how learnable multi-head self-attention (MHSA) modules can perform over regular self-attention used in ViTs, a defining technology in vision in the last six years. Taken together, our contributions are:
Designing a Learnable Attention Module for ViTs. This pilot study is on the self-attention in ViTs. In §3, we propose a general learnable Kolmogorov-Arnold multi-head self-attention, or KArAt, for ViTs that can operate on any basis. Due to the memory explosion, any full-rank learnable MHSA operator in ideal KArAt causes a training bottleneck. We propose a low-rank approximation of the operator to reduce memory requirements to a feasible level, but leave the possibility of exploring alternative techniques. By adopting the Fourier basis into this modular version, we design Fourier KArAt; see §4 for details. We cannot perform tasks with varying sequences and cross-attention without major modifications to KArAt. The most prominent modalities other than vision, such as audio, speech, video, and text, have varying sequence length, involve modeling with cross-attention-based encoder-decoder architectures, for which token interactions would be substantially different from self-attention. These are non-trivial technical challenges for implementing KArAt.
Benchmarking, Evaluation, and Analysis (§56). We benchmark Fourier KArAt on CIFAR-10, CIFAR-100, krizhevsky2009cifar10/100 and ImageNet-1K deng2009imagenet datasets and evaluate their performance against vanilla ViTs (ViT-Tiny, -Small wu2022tinyvit ; steiner2022howtotrainvit , and -Base dosovitskiy2020vit ), and other popular ViTs (ConViT convit and Swin-Transformer liu2021swin ). We use small-scale datasets, SVHN svhn , Oxford Flowers 102 flowers102 , and STL-10 stl10 for understanding the transfer learning capability of the models. Additionally, we embed KArAt with ViT-Det vitdet for the object detection and segmentation on MS COCO coco . We dissect KArAts’ performance and generalization capacity by analyzing their loss landscapes, weight distributions, optimizer path, and attention visualization, and compare them with traditional softmax attention in vanilla ViTs. In addition, we investigate KArAts’ explanability for visual understanding; see §5.3.
2 Background
Multi-layer Perceptrons (MLPs) consist of $L$ layers. By convention, $a^{[0]}=x^{[0]}$ denotes the input data and the $l^{\rm th}$-layer is defined for input $x^{[l-1]}\in\mathbb{R}^{d_{l-1}}$ as
$a^{[l]}=\Phi^{[l]}(W_{l}x^{[l-1]}+b_{l}),\ W_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}},b_{l}\in\mathbb{R}^{d_{l}}.$}
In MLP, the nonlinear activation functions, $\Phi(\cdot)$, are fixed bengio2017deep . For supervised tasks, given a training dataset $D$ with $N$ elements of the form (input, ground-truth) pairs, $\{(x_{i},y^{\star}_{i})\}_{i=1}^{N}$, the loss, $\mathcal{L}(X,Y^{\star}|\mathcal{W})=d_{Y}(a^{[L]},Y^{\star})$ (metric induced by the space $\mathbb{R}^{d_{L}}$) is calculated in the forward pass. The layerwise weights, $\{(W_{l},b_{l})\}_{l\in[L]}$ are learned by minimizing $\mathcal{L}(X,Y^{\star}|\mathcal{W})$.
Kolmogorov-Arnold Network (KAN) liu2024kan is a neural network involving learnable activations parametrized by a chosen set of basis functions defined on a set of grid points or knots. The idea for this network stems from the Kolmogorov-Arnold Representation Theorem.
Theorem 1 (Kolmogorov-Arnold Representation Theorem).
kolmogorov1956 For any multivariate continuous function, $f:[0,1]^{n}\to\mathbb{R}$, there exists a finite composition of continuous single-variable functions, $\phi_{q,p}:[0,1]\to\mathbb{R},\Phi_{q}:\mathbb{R}\to\mathbb{R}$ such that $f(x)=f(x_{1},x_{2},\cdots,x_{n})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right).$
Theorem 1 describes an exact finite expression using 2 layers, but liu2024kan generalized this representation to multi-layers via KAN; it is analogous to the Universal Approximation Theorem hornik1989multilayer . For input $x^{[0]}$, an $L$-layer KAN is given as $\text{KAN}(x^{[0]})=\Phi^{L}\circ\cdots\circ\Phi^{l}\circ\cdots\circ\Phi^{1}(x^{[0]})$. In each KAN layer, the activation function $\Phi$ is learnable, and $\Phi^{[l]}=[\Phi_{ij}^{[l]}]$} operates on $x^{[l-1]}\in\mathbb{R}^{d_{l-1}}$} to produce $x^{[l]}\in\mathbb{R}^{d_{l}}$}, such that:
${x^{[l]}_{i}=\Phi^{[l]}_{i:}(x^{[l-1]})=\sum_{j=1}^{d_{l-1}}\Phi^{[l]}_{ij}(x^{[l-1]}_{j})}$.
Originally, liu2024kan chose the B-spline basis functions. That is, the activation functions were defined to be $\phi(x)=w(b(x)+\text{spline}(x))$, where $b(x)=\texttt{SiLU}(x)=\frac{x}{1+e^{-x}}$ and $\text{spline}(x)=\sum\limits_{i}c_{i}B_{i}(x)$, with $B_{i}(\cdot)$ being one of the $k$-th degree B-spline basis functions parametrized by the $G$ grid points on a uniform grid, [I,I][-I,I]. In this case, the representation for each function is given as:
$).\Phi_{ij}^{[l]}(x^{[l-1]}_{j})=w^{[l]}_{ij}(\texttt{SiLU}(x^{[l-1]}_{j})+\sum\limits_{m=1}^{G+k-1}c^{[l]}_{ijm}B_{ijm}(x^{[l-1]}_{j}$)).
The weights $\{([c^{[l]}_{ijm}],[w^{[l]}_{ij}])\}_{l=1}^{L}$} are learned by minimizing the loss for a supervised learning task. The bases for KANs’ activations could be Fourier mehrabian2024implicit ; xu2024fourierkan , wavelet bozorgasl2024wav , fractals yang2024kolmogorov , etc.
Multi-head self-attention (MHSA) in ViTs dosovitskiy2020vit . The encoder-only vanilla ViT architecture is inspired by the Transformer proposed in attention . For simplicity, we do not mention the details of the layer normalization and other technicalities. Our central focus is the MHSA architecture. For limited space, we move this discussion to §A.
3 How Can We Design Learnable Attention?
Last year, we witnessed a surge in embedding KANs in different DNN architectures genet2024tkan ; ferdaus2024kanice ; li2024UKAN . To our knowledge, the work closely related to ours is KAT, where yang2024kolmogorov replaces the MLP layers in the ViTs with KAN layers; also, see VisionKAN2024 . We note that, for attentive graph neural networks (GNNs), fang2025kaa unifies the scoring functions of attentive GNNs and names it as Kolmogorov-Arnold Attention (KAA). However, it is orthogonal to our MHSA design; deploying learnable MHSA is complicated. In this Section, we design learnable multi-head self-attention (MHSA) module for ViTs, with improved interpretability, adaptability, and expressiveness that operate inside its encoder blocks.
Let ${\cal A}^{i,j}\in\mathbb{R}^{N\times N}$ be the attention matrix for $i^{\rm th}$ head in the $j^{\rm th}$ encoder block. Instead of using the softmax function row-wise, we can use a learnable activation function, $\tilde{\sigma}$} on the row vectors of each attention head ${\cal A}^{i,j}$. With any choice of the basis functions (e.g., B-Spline, Fourier, Wavelets, etc.), the activated attention row vector, $\tilde{\sigma}(\mathcal{A}^{i,j}_{k,:})$} for $k\in[N]$ can be written as
$\textstyle{\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]}$ $=\textstyle{\begin{pmatrix}\phi_{11}(\cdot)&\phi_{12}(\cdot)&...&\phi_{1N}(\cdot)\\ \phi_{21}(\cdot)&\phi_{22}(\cdot)&...&\phi_{2N}(\cdot)\\ \vdots&\vdots&\ddots&\vdots\\ \phi_{N1}(\cdot)&\phi_{N2}(\cdot)&...&\phi_{NN}(\cdot)\end{pmatrix}}\scriptscriptstyle{\begin{pmatrix}{\cal A}^{i,j}_{k,1}\\ {\cal A}^{i,j}_{k,2}\\ \vdots\\ {\cal A}^{i,j}_{k,N}\end{pmatrix}}=\textstyle{\left(\Phi^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\right)^{\top}},$ (1)
where $\Phi^{i,j}=[\phi^{i,j}_{pq}]\in\mathbb{R}^{N\times N}$ and each matrix entry, $\phi_{pq}$, is referred to as a learnable unit, represented using a set of basis functions $\{\psi^{i,j}_{m}\}_{m=1}.$ The coefficients associated with the basis functions are the learnable parameters. In our convention, ${\cal A}^{i,j}_{k,:}\in\mathbb{R}^{1\times N}$} is a row vector. We apply the transpose operation to each row vector to adopt the convention used in KAN layers. Hence $\Phi^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\in\mathbb{R}^{N\times 1},$} and we transpose it to obtain learnable attention row vectors $\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]$}; see Figures 1(b)(ii)-(iii).
Projection onto the probability simplex. The softmax function acts on each row of $\cal A$} to create a probability distribution; the learnable attention does not guarantee that. To ensure that each row vector of $\Phi(\mathcal{A})$ lies on a probability simplex, we project them onto the $\ell_{1}$-unit ball. That is, for each attention matrix, ${\cal A}^{i,j}\in\mathbb{R}^{N\times N}$, and for each $k\in[N],$ we want to have $\|\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]\|_{1}=1$} and $\tilde{\sigma}\left[({\cal A}^{i,j}_{k,l})\right]\geq 0$ for $l\in[N]$. We cast a sparse approximation problem, whose variants have been well-studied in the past decade and arise frequently in signal processing bryan2013making ; donoho2006compressed ; dutta_thesis and matrix approximation boas2017shrinkage ; stewart1993early ; see §3. Algorithm 1 provides an $\ell_{1}$-projection pseudocode; for each $k\in[N]$, setting $y=\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]$} and $m=N$ in Algorithm 1, we obtain the projected attention vector.
4 Our Architecture
Different from learning activations for MLP, as in KAN and KAT, implementing the generic learnable attention incurs impractical computational and memory costs on present computing hardware and DL toolkits. For learnable activation, $\tilde{\sigma}:\mathbb{R}^{N}\to\mathbb{R}^{N}$, a full-rank operator, $\Phi^{i,j}\in\mathbb{R}^{N\times N}$} acts on each row ${\cal A}^{i,j}_{k,:}$ of ${\cal A}^{i,j}$, and produces $\tilde{\sigma}({\cal A}^{i,j}_{k,:}).$ These operations are extremely compute-heavy and have large memory footprints. E.g., ViT-Tiny dosovitskiy2020vit ; wu2022tinyvit ; steiner2022howtotrainvit has 5.53M parameters and requires nearly 0.9 GB of GPU memory to train. Implementing the learnable attention equation 1 in ViT-Tiny training with B-splines of order 5 and grid size 10 and Fourier basis with grid size 10 increases the parameter count to 30.68M and 39.06M, respectively. We could not train the version with B-splines for its humongous memory requirements, and B-spline computations are non-parallelizable. The one with the Fourier basis is parallelizable, but it takes approximately 60 GB of GPU memory when computing with a batch of one image. This computational bottleneck is agnostic of the basis functions.
Figure 2: Different configurations to update $\widehat{\Phi}$:(a) Blockwise configuration, where $\Phi^{i,1}\neq\Phi^{i,2}\neq\cdots\neq\Phi^{i,L}$ for all i=1,2,,hi=1,2,...,h; (b) universal configuration, where $\Phi^{i,1}=\Phi^{i,2}=\cdots=\Phi^{i,L}=\Phi^{i}$ for all i=1,2,,h.i=1,2,...,h.
How can we remedy this? Deep neural networks exhibit low-rank structure oja1982simplified ; jain2016streaming . Recently, kwon2024efficient showed that across various networks, the training largely occurs within a low-dimensional subspace. Therefore, we postulate that the attention matrices also have an underlying low-rank structure.
To validate this, we perform spectral analysis of the attention matrices before and after the softmax activation; see Figures 5-5. The scree test cattell1966scree shows that for traditional attention matrices without softmax, there are 8 significant singular values; after softmax activation, this count increases to 16. The above observation verifies that with or without the nonlinear activation, the attention matrix $\mathcal{A}^{i,j}$ possesses an underlying low-rank structure. This motivates us to use a lower-dimensional operator, $^{\widehat{\Phi}}$ for the learnable attention calculation with $r=12$. In §C.4.2-C.4.3, we perform an ablation to find the best $r.$
Instead of $\Phi^{i,j}\in\mathbb{R}^{N\times N}$}, we use an operator, $\widehat{\Phi}^{i,j}\in\mathbb{R}^{r\times N}$} such that $r\ll N$}, and the new learned activation is $r$-dimensional for $k\in[N]$.  This process significantly reduces the computational overhead. Next, we post-multiply another learnable weight matrix, $W^{i,j}\in\mathbb{R}^{N\times r}$ to project them back to their original dimension. For each $k\in[N]$}, this operation results in computing $\widehat{\sigma}({\cal A}^{i,j}_{k,:})=\left[W^{i,j}\widehat{\Phi}^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\right]^{\top}$} (Figure 1(b)(iv)); see other variants in §C.5. Our approximation is not unique or the best; it is one of the feasible solutions given the present computing interface.
Fourier Kolmogorov-Arnold Attention. Although the default basis for KANs is B-Splines in liu2024kan , the number of MHSA parameters is much more than that of an MLP. Specifically, for $L$ encoder blocks, each with $h$ attention heads, the parameter complexity is $O(N^{2}Lh(G+k))$, for $\Phi\in\mathbb{R}^{N\times N}$ and $O(NrLh(G+k))$, for $\widehat{\Phi}\in\mathbb{R}^{r\times N}$, if the model uses B-Splines of degree $k$.
Moreover, as yang2024kolmogorov mentioned, B-splines are localized functions, not standard CUDA functions. Although efficient CUDA implementations for cubic B-Splines exist ruijters2012gpu ; ruijters2008efficient , their overall implementation results in slower, sparse, non-scalable, and complicated GPU execution; also, see xu2024fourierkan ; KANsLimits . In §C.4.1, our extensive experiments with B-spline basis in KAN for image classification tasks verify their poor generalizability on medium-scale datasets (e.g., CIFAR-10 and CIFAR-100).
So, what could be an attractive basis? A fundamental question in function approximation is whether the function converges pointwise almost everywhere. Carleson in 1966 proved the following fundamental result for the Fourier approximation of $L^{p}$ periodic functions.
Theorem 2.
carleson1967convergence Let $f$ be an $L^{p}$} periodic function for $p\in(1,\infty]$}, with Fourier coefficients $\hat{f}(n)$}. Then $\lim_{N\to\infty}\sum_{|n|\leq N}{\hat{f}}(n)e^{inx}=f(x)$}, for almost every $x$.
In particular, if a function is continuously differentiable, its Fourier series converges to it everywhere. In the past, xu2024fourierkan ; mehrabian2024implicit used the Fourier basis in KAN, dong2024fan proposed a Fourier analysis network (FAN) to capture periodic phenomena. Motivated by these, we use the Fourier basis to approximate the effect of the smooth softmax function. We employ the Fourier basis, $\{\sin{(\cdot)},\cos{(\cdot)}\}$ with gridsize $G$ to design learnable attention. Algorithm 1 can be used optionally. For $\widehat{\Phi}^{i,j}=[\widehat{\phi}^{i,j}_{pq}]\in\mathbb{R}^{r\times N}$}, we have $\widehat{\Phi}^{i,j}:\mathbb{R}^{N}\to\mathbb{R}^{r}$}. Hence, each row ${\cal A}^{i,j}_{k,:}$}, for $k\in[N]$}, transformed into $\widehat{\Phi}^{i,j}_{p}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]=\sum_{q=1}^{N}\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})$}, via the Fourier bases such that
$\textstyle{\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})=\sum_{m=1}^{G}a_{pqm}\cos{(m{\cal A}^{i,j}_{k,q})}+b_{pqm}\sin{(m{\cal A}^{i,j}_{k,q})},}$ (2)
where $m$ refers to the $m^{\rm th}$} grid point on a uniform grid of size $G$. The weights, $\{[a_{pqm}],[b_{pqm}]\}_{m=1}^{G},W^{i,j}]$} are updated in the backpropagation. If we use a Fourier basis in $\widehat{\Phi}\in\mathbb{R}^{r\times N}$}, the parameter complexity of MHSA becomes O(2NrhGL).{O(2NrhGL)}. Regardless of the basis used, the linear projection takes $O(N^{2}rhL)$ FLOPs. For any Transformer, $L$ and $h$ are fixed. So, one may omit them from the complexity results. See FLOPs and memory requirement in Table 13.
Table 1: Performance of the best-performing Fourier KArAt models compared to the conventional vanilla ViT baselines. The best and the second-best Top-1 accuracies are given in red and blue, respectively. The $\downarrow$ and $\uparrow$ arrows indicate the relative loss and gain, respectively, compared to the base models.
Model CIFAR-10 CIFAR-100 ImageNet-1K Parameters
Acc.@1 Acc.@5 Acc.@1 Acc.@5 Acc.@1 Acc.@5
ViT-Base 83.45 99.19 58.07 83.70 72.90 90.56 85.81M
[…]
ViT-Small 81.08 99.02 53.47 82.52 70.50 89.34 22.05M
[…]
ViT-Tiny 72.76 98.14 43.53 75.00 59.15 82.07 5.53M
[…]
Table 2: Performance of Fourier KArAt fine-tuned on small datasets from ImageNet-1K pre-trained weights.
Model Acc.@1
SVHN Flowers 102 STL-10
ViT-Base 97.74 92.24 97.26
+ $G_{1}B$ 96.83 89.66 95.30
+ $G_{1}U$ 97.21 89.43 95.78
ViT-Small 97.48 91.46 96.09
+ $G_{3}B$ 97.04 89.67 95.26
+ $G_{3}U$ 97.11 90.08 95.45
ViT-Tiny 96.69 84.21 93.20
+ $G_{3}B$ 96.37 83.67 93.09
+ $G_{3}U$ 96.39 83.70 92.93
Blockwise and Universal operator configuration. We consider two configurations for updating the operator $\widehat{\Phi}^{i,j}.$ (a) Blockwise: In this configuration, each encoder block learns the attention through $h$ distinct operators $\widehat{\Phi}$ for each of the $h$ heads, totaling $hL$ operators; see Figure 2 (a). Like the MHSA in vanilla ViTs, the blockwise configuration is designed to learn as many different data representations as possible. (b) Universal: This is motivated from the KAT yang2024kolmogorov . In KAT, the MLP head is replaced with different variations of a KAN head—KAN and Group-KAN. In Group-KAN, the KAN layer shares rational base functions and their coefficients among the edges. Inspired by this, in our update configuration, all $L$ encoder blocks share the same $h$ operators; $\widehat{\Phi}^{i,j}=\widehat{\Phi}^{i}$ for $j=1,2,\ldots,L$; see Figure 2-(b). Rather than learning attention through $hL$ operators, this configuration only uses $h$ operators. Here, we share all learnable units and their parameters across $L$ blocks in each head. We postulate that the blockwise mode with more operators captures more nuances from the data. In contrast, the universal mode is suitable for learning simpler decision boundaries from the data. We also note that these two modes can be used with equation 1 as shown in Figure 1(b). Finally, Algorithm 2 in the Appendix provides the pseudocode for applying Fourier-KARAT in the ViT encoder block; see more nuanced discussion about them in §C.4.6.
5 Empirical Study
Implementing KArAt. We provide the implementation details in §C.1; see §C.1.1 for baselines and datasets. We integrate 5 wavelet bases, rational basis, Fourier basis, and 3 different base activation functions in our learnable MHSA design; see §C.1.3. Fourier KArAt performs the best, and we report it in the main paper. It has 1 hyperparameter, the grid size, $G$, and 2 configurations, blockwise or universal. We vary grid sizes from the set $G=\{1,3,5,7,9,11\}.$} For each value, we perform universal and blockwise updates; see Figure 2. With this formalization, $G_{n}B$ and $G_{n}U$ denote Fourier-KArAt with grid size $n$ and weights updated in blockwise and universal mode, respectively. See the ablation study of these variants in §C.5. We dispense the $\ell_{1}$ projection, as projecting the learned attention vectors to a probability simplex degrades the performance; see §C.4.5.
5.1 Trained Model Quality, Transferability, and Generalizability
Image Classification. Table 2 presents the performance of the best-performing Fourier KArAt variants; see Figure 6 for the training loss and test accuracy curves. Fourier KArAt is easily optimized in the initial training phase, gaining at par or better loss value and accuracy than the vanilla ViTs. However, in the later training phase, except for the fewer parameter variants (ViT-Tiny+Fourier KArAt), the loss curve flattens faster; we infer that for most of the Fourier-KArAt scenarios. We also notice that larger models like ViT-Base are more susceptible to slight changes in the parameters, making the training process challenging to manage, which we try to investigate in the rest of the paper using several analytical studies. Overall, ViT-Tiny+Fourier KArAt variants outperform ViT-Tiny on CIFAR-10 and CIFAR-100 by $5.40\%$ and $7.40\%$, respectively, and on ImageNet-1K, it achieves a similar accuracy. ViT-Small and -Base models with Fourier KArAt variants can barely outperform the vanilla models, and the accuracy differences increase as we tackle larger base models.
Performance on other ViTs for Image Classification. We fused KArAt with modern ViT-based architectures, ConViT convit and Swin Transformer liu2021swin ; §C.1.4 provides their implementation details. Table 4 shows that Fourier KArAt outperforms ConViT on CIFAR-10, experiences a modest drop on ImageNet-1K, and shows lower performance for hierarchical models, such as Swin, which requires a careful investigation.
Transferability: Image Classification. Unlike ViTs, which were pre-trained on large datasets and then transferred to various mid-sized or small image recognition benchmarks dosovitskiy2020vit , which helped them achieve state-of-the-art results, KArAts were never pre-trained with such datasets to showcase their full potential. Due to limited computing resources, we could not perform large-scale training. We investigate the KArAts’ transfer capability by fine-tuning on smaller datasets like SVHN svhn , Oxford Flowers 102 flowers102 , and STL-10 stl10 from their ImageNet-1K pre-trained weights; see dataset details in §C.1.1. Table 2 shows KArAts transfer well across all datasets and have comparable performance with vanilla ViTs, even when their ImageNet-1K performance was not equivalent.
Generalizability: Detection & Segmentation. To understand the robustness of the Fourier KArAt, we employ ViT backbones for detection and segmentation; see results and discussion in §C.2.
Key Takeaways. (a) The Top-1 accuracies from Tables 2 and 2 show Blockwise configuration is more desirable over Universal for image classification. (b) From Table 2, we can infer that the Fourier KArAt transfers well in fine-tuning tasks. However, the transferability of hyperparameters, e.g., grid size $G$, across datasets remains an open question. (c) KArAt generalizes well. For random initialization in detection, the vanilla ViT-Det has an advantage over KArAt, and thus, we see a small performance gap. The best KArAt hyperparameters for this task are yet to be found. Overall, KArAt shows significant potential if the incompatibilities are properly addressed.
Table 3: Performance comparison of Fourier-KArAt and softmax attention on various vision transformer architectures.
Model CIFAR-10 ImageNet-1K
Acc.@1 Acc.@5 Acc.@1 Acc.@5
ConViT-Tiny 71.36 97.86 57.91 81.79
+ $G_{3}B$ 75.57 98.61 56.57 80.75
+ $G_{3}U$ 74.51 98.63 56.51 80.93
Swin-Tiny 84.83 99.43 76.14 92.81
+ $G_{3}B$ 79.34 98.81 73.19 90.97
Table 4: Attention Transfer from ViT+Fourier-KArAt to vanilla ViTs, following li2024surprising .
Model Attention CIFAR-10 CIFAR-100
Transfered From Acc.@1 Acc.@5 Acc.@1 Acc.@5
ViT-Tiny None 72.76 98.14 43.53 75.00
[…]
ViT-Base None 83.45 99.19 58.07 83.70
[…]
5.2 Performance Analysis
We dissect the performance of Fourier KArAt using the following tools:
[…]
Attention Visualization. MHSA in vanilla ViTs captures region-to-region interaction in images. DINO caron2021dino provides an innovative way to explain the dominant regions in an image that contribute towards the inference decision. It maps the self-attention values of the CLS token at the last encoder layer, which is directly responsible for capturing the information related to the class label. While the computer vision community predominantly uses this attention visualization, it is unsuitable for our case. Unlike softmax, Fourier KArAt does not inherently ensure learned attention values in $[0,1]$, as we dispense the $\ell_{1}$ projection. Therefore, we cannot interpret its performance similarly. Nonetheless, pre-softmax or pre-KArAt values in the attention matrix $\mathcal{A}^{i,j}$ are supposed to capture token-to-token interactions, and we adapt the attention maps to ignore the negative values for the sake of visualization. We use the ImageNet-1K trained models of ViT-Tiny (with traditional MHSA and Fourier KArAt) and plot the attention maps of the last layers in Figures 4 and 11.
Figure 3: 3D-visualization of Loss landscape for ViT-Tiny and ViT-Base along the two largest principal component directions of the successive change of model parameters. KArAt’s loss landscapes are significantly less smooth than those of traditional attention; spiky loss landscapes are undesirable for optimization stability and the generalizability of the resulting model. See Figure 10 for the loss contours and the optimizer trajectory.
Figure 4: Vit-Tiny Attention map visualization. Original images for inference (the left), the attention score (middle), and image regions of the dominant head (Top row: Fourier KArAt, bottom row: traditional MHSA).
5.3 KArAt’s Flexibility over Traditional Attention in ViT
Stemming from Kolmogorov-Arnold Representation Theorem (Theorem 1kolmogorov1956 , KANs liu2024kan with learnable activation functions can approximate complex mathematical functions and symbolic representations, and provide an interpretable alternative to MLPs liu2024kan ; liu2024kan2 ; yang2024kolmogorov . Precisely, if the target function admits a structure of the composition of $L$-KAN layers with smooth activation functions, then one can use an $L$-layer KAN with B-Spline basis to approximate it well [Theorem 2.1 in liu2024kan ]. This allows them to model complex relations that lie in the data distribution; KANs can discover the need for a new function whose numerical behavior suggests it may be a Bessel function (Figure 23(d) in liu2024kan ), and the authors also show that KANs can discover unknown equations (Figure 4(e)).
By using a similar argument, we hypothesize that KANs could be a better fit for approximating the underlying non-linear relationship between image patches and can be used for modeling token interactions with learnable components. Hence, KArAt can be viewed as an upgrade to vanilla self-attention, offering adaptability and more freedom for modeling token interactions in an interpretable way. KArAt provides the freedom to choose from a multitude of simple, easy-to-interpret, and tunable functions whose linear combination can produce a large class of unknown functions, rather than a fixed and known function, softmax. The traditional softmax attention adds a non-linearity for modeling such interactions in a one-size-fits-all mechanism without any justification. This interpretability is directly reflected in KArAt’s attention maps; see Figures 4 and 11.
Why KArAt’s attention maps are more explainable? DINO caron2021dino , being self-supervised, attends to the entire object and obtains richer features and deep visual representations ericsson2021well . To attend to the entire object like self-supervision, chefer2022optimizing in their supervised training, optimizes the attention map with additional loss objectives, which improves model interpretability and explanability. KArAt’s attention maps in Figures 4 and 11 capture entire objects without any extra loss, unlike some contours in the traditional softmax attention. It provides improved explanability and visual understanding, compared to spurious cues in supervised ViTs.
To show that KArAt’s attention maps are not just visually better, but the learnable attention leads to effective token interactions modeling, we follow the study by li2024surprising . To understand how the attention modulates the performance of ViTs, li2024surprising considered attention matrices from a masked auto-encoder he2022masked , trained with self-supervision, having superior performance, and used them in a vanilla ViT, ignoring the query and key, and training the rest of the network. This study shows that attention is the most important component in a transformer and can bring improved performance over the vanilla supervised ViT. In the same spirit, we use the precomputed attention matrices from the KArAt models to train the value branch of a vanilla ViT from scratch; key and query branches remain frozen. See the implementation details in §C.1 and the results in Table 4.
[…]
6 Discussion and Conclusion
The learnable MHSA design raises many interesting questions regarding KArAt’s computing scalability and the resulting models’ generalizability. Below, we discuss their potential positive (++) and negative (-) implications and encourage researchers to elaborate on them.
Designing KArAt self-attention is not parameter efficient and they are memory hungry (-). Having learnable activations in the MHSA causes a memory explosion on the present GPUs. Hence, calculations with the full-rank operator, $\Phi\in\mathbb{R}^{N\times N}$}, are prohibitively expensive, regardless of the basis. Present GPUs and CUDA functions are not optimized for KAN implementation as noted by many before us; see yang2024kolmogorov ; ruijters2012gpu ; ruijters2008efficient ; xu2024fourierkan ; KANsLimits . Due to this, although parameter increase is not always significant (Table 2), optimized Fourier KArAt variants utilize $2.5-3\times$ more GPU memory compared to traditional softmax attention. We propose a low-rank approximation to reduce memory requirements to a feasible level (Table 13), but researchers should explore alternative techniques. Interestingly, while there is a significant training time discrepancy between vanilla ViTs and Fourier-KArAts, their inference speeds are comparable on the present compute interface and hardware configuration; see Figure 7 in §C.3. Hence, we hope the future holds more efficient system-level optimization recipes for training KArAt. Additionally, we witnessed KArAts’ inconsistent training behavior with parameter changes in §C.4.4. Each ViT model with particular Fourier-KArAt variants has a typical $G$ value that brings out its best performance; there is no universal $G$ to follow.
KArAts show a better attention score (++) and transfer learning capability (++) for all ViTs, but their generalizability does not scale with larger ViTs (-). Fourier KArAt concentrates high interaction scores on the entire object compared to spurious cues chefer2022optimizing like vanilla ViTs dosovitskiy2020vit ; deit . They exhibit decent transfer learning and generalizability. The distribution of weights guarantees the stability of the training. Hence, by zhang2022neural , we can postulate that first-order optimization algorithms (e.g., ADAM) can optimize KArAt’s loss landscape. But zhang2022neural cannot explain why KArAt’s generalizability does not scale with larger models due to high parameterization. We encourage researchers to investigate these models’ local or global optimality and generalizability through theoretical and empirical studies.
Overparameterized KArAts’ weights lie in a much lower-dimensional subspace (++). Spiky loss landscape with local minima is probably the major impediment towards KArAt’s scalability; see Figure 3. Modern transformers are highly over-parameterized, and their weight matrices show low-rank structures kwon2024efficient ; spectral analysis (§C.7) of KArAt shows its attention matrices possess a better low-rankness than traditional attention (Figure 13). Considering this, properly guided parameter reduction in KArAt can result in an optimized model that generalizes well at scale.
KArAts can provide better interpretability for smaller models (++). KANs are claimed to be more interpretable and accurate than MLPs due to their flexible choice of univariate functions liu2024kan . It is evident from KArAts’ attention maps, which are directly responsible for performance improvement; see §5.3. ViT-Tiny+KArAts outperforms vanilla ViT-Tiny on small-scale datasets, and performs at par on ImageNet-1K (Table 2) and transfer learning (Table 2). In the future, KArAt’s learnable activation makes a case for investigating it to add more explanability in vision tasks and can find more interpretability in smaller models trained in a limited data scenario.
Conclusion.  Across our benchmarking, KArAt outperforms traditional MHSA in smaller ViTs; results on larger models are mixed. Additionally, we note that the computing requirement impedes KArAt’s performance—training and fine-tuning these models are expensive. However, learnable activations can make a substantial difference in a model’s interpretability, and these limitations are not KArAt’s weaknesses. At its early stage, KArAt’s decent performance on diverse tasks shows its potential. In the future, one can extend this study beyond ViTs and check the resilience of KArAt in language processing and on large multimodal models llama3 ; qwen2.5 ; thawakar2024mobillama ; campos2025gaea ; gpt3 ; llama32vision ; team2023gemini ; flamingo .
Appendix
Organization. We organized the Appendix as follows: §B contains the solution to problem equation 3. We also provide the pseudocode to implement Fourier-KArAt in ViT’s encoder block; see Algorithm 2. §C discusses additional numerical results, including implementation details, hyperparameter tuning, computation time analysis, diverse ablation studies, and different variants of Fourier KArAt. Finally, we concluded §C with the spectral analysis of the attention matrices for traditional and learnable attention.
Appendix A Multihead Self Attention—Continued
Let $X\in\mathbb{R}^{H\times W\times C}$} be an input image of resolution ${H\times W}$. ViT uses a $p\times p$ patch and generates $N=\tfrac{HW}{p^{2}}$ input tokens. Each token is vectorized along $C$ channels to produce $X_{T}\in\mathbb{R}^{N\times p^{2}C}$}, where each row, ${X_{T}}_{i,:}$, represents a flattened patch across channels. A learnable embedding matrix, $E\in\mathbb{R}^{p^{2}C\times d}$ projects ${X_{T}}_{i,:}$ to an embedding vector in $\mathbb{R}^{d}$ such that after appending the learnable class token, $x_{{\cal G}}$ and adding the positional encoding matrix, $P_{E}\in\mathbb{R}^{(N+1)\times d}$ of compatible size, the final input, $X_{I}=[x_{{\cal G}};X_{E}]+P_{E}$. The matrix $X_{I}$ is layer normalized and further projected on the row spaces of three learnable weight matrices, $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$ to generate the query, key, and value matrices, $Q,K,$ and $V$, respectively, via $Q=X_{I}W_{Q},K=X_{I}W_{K},\;{\rm and}\;V=X_{I}W_{V}.$ These matrices are further divided into $h$ partitions (also, called heads), $Q^{i},K^{i},V^{i}\in\mathbb{R}^{N\times\tfrac{d}{h}}$}, such that each partition generates a self-attention matrix for the $i^{\rm th}$ head, ${\cal A}^{i}=\tfrac{Q^{i}{K^{i}}^{\top}}{\sqrt{d/h}}.$} Finally, we project $V^{i}$ onto the column space of $\sigma({\cal A}^{i})$ as $\sigma({\cal A}^{i})V^{i}$, where $\sigma$ is the softmax operator applied row-wise. The outputs of different heads are further concatenated into a large matrix $\begin{bmatrix}\sigma({\cal A}^{1})V^{1}&\sigma({\cal A}^{2})V^{2}&\cdots&\sigma({\cal A}^{h})V^{h}\end{bmatrix}W^{O},$} post-multiplied by the learnable weight, $W^{O}\in\mathbb{R}^{d\times d}$. This process is sequentially performed in $L$ encoder blocks; see Figure 1(b)(i).
Appendix B A Sparse Approximation Problem and its Solution
Let $y\in\mathbb{R}^{m}$ be a given vector. If we want to approximate $y$ with a vector, $x\in\mathbb{R}^{m}$, with positive components and $\|x\|_{1}=1$, we can write the constrained optimization problem as:
[…]
First, we rewrite the constrained problem equation 3 as an unconstrained problem using the Lagrange multipliers as:
${\cal L}(x,\lambda,\mu)=\frac{1}{2}\|x-y\|^{2}+\lambda\bigg(\sum\limits_{i=1}^{m}x_{i}-1\bigg)-\mu^{T}x,$ (4)
where $\lambda,\mu$ are Lagrange multipliers. Using the Karush-Kuhn-Tucker (KKT) stationarity condition on equation 4, we find $0\in\partial{\cal L}(x,\lambda,\mu)$, which implies, $x_{i}-y_{i}+\lambda=\mu_{i}.$ The complementary slackness gives, $\mu_{i}x_{i}=0$ for $i\in[m].$ Further, for $i\in[m]$, the primal and dual feasibility conditions are $x_{i}\geq 0,$ $\sum\limits_{i=1}^{m}x_{i}=1$, and $\mu_{i}\geq 0,$ respectively.
The stationarity and complementary slackness conditions give, $x_{i}-y_{i}+\lambda\geq 0,\ (x_{i}-y_{i}+\lambda)x_{i}=0$. If $\lambda\geq y_{i}$, then $x_{i}=0$. Otherwise, we have that $\sum\limits_{i=1}^{m}x_{i}=\sum\limits_{i=1}^{m}(y_{i}-\lambda)=1$ implying $\lambda=\frac{1}{m}(\sum\limits_{i=1}^{m}y_{i}-1)$.
Algorithm 1 Projection on $\ell_{1}$ ball [83]
Input: $y\in\mathbb{R}^{m}$
Sort: $y$ as $y_{(1)}\geq y_{(2)}\geq...\geq y_{(m)}$
Calculate: $\rho\triangleq\max\limits\Big\{i\in[m]|\ y_{(i)}-\frac{1}{i}\bigg(\sum\limits_{j=1}^{i}y_{(j)}-1\bigg)\Big\}$
Set: $\lambda=\frac{1}{\rho}\bigg(\sum\limits_{i=1}^{\rho}y_{(i)}-1\bigg)$
Output: $x^{\star}=(y-\lambda)_{+}\triangleq\max{(y-\lambda,0)}$
Algorithm 2 Fourier-KArAt in $j^{\rm th}$ Vision Transformer block
Input: $X\in\mathbb{R}^{N\times d},\text{ Fourier basis }\{\sin{(\cdot)},\cos{(\cdot)}\}$
Output: $O\in\mathbb{R}^{N\times d}$
Parameters: $W_{Q}^{i},W_{K}^{i},W_{V}^{i}\in\mathbb{R}^{d\times d_{h}},d_{h}=\frac{d}{h},G,[\{[a_{pqm}],[b_{pqm}]\}_{m=1}^{G},W^{i,j}],W^{O,j}$
Hyperparameters: Blockwise, universal, Algorithm 1 ($\ell_{1}$ projection), $L$ encoder blocks
for each head $i\in\{1,\ldots,h\}$ do:
     $Q^{i,j}\leftarrow XW^{i,j}_{Q}$
     $K^{i,j}\leftarrow XW^{i,j}_{K}$
     $V^{i,j}\leftarrow XW^{i,j}_{V}$
     ${\cal A}^{i,j}=\frac{Q^{i,j}(K^{i,j})^{\top}}{\sqrt{d_{h}}}$
end for
for each head $i\in\{1,\ldots,h\}$ do:
     for each row $k\in\{1,\ldots,N\}$ do:
      $\widehat{\Phi}^{i,j}[({\cal A}^{i,j})_{k,:}^{\top}]\leftarrow\sum_{q=1}^{N}\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})=\sum_{m=1}^{G}a_{pqm}\cos{(m{\cal A}^{i,j}_{k,q})}+b_{pqm}\sin{(m{\cal A}^{i,j}_{k,q})}$
      if $\ell_{1}$ projection then
       Execute Algorithm 1
      else
       pass
      $\widehat{\sigma}({\cal A}^{i,j}_{k,:})\leftarrow\left[W^{i,j}\widehat{\Phi}^{i,j}[({\cal A}^{i,j}_{k,:})^{\top}]\right]^{\top}$
     end for
end for
$O\leftarrow\begin{bmatrix}\widehat{\sigma}({\cal A}^{1,j})V^{1,j}&\widehat{\sigma}({\cal A}^{2,j})V^{2,j}&\cdots&\widehat{\sigma}({\cal A}^{N,j})V^{N,j}\end{bmatrix}W^{O,j}$
if Blockwise Mode then
     Return $O$
else if universal Mode then
     pass learnable units and parameters to $(j+1)^{\rm th}$ encoder block
     Return $O$
Note that, in Algorithm 1, one can replace $y$ with $\widehat{\Phi}^{i,j}[({\cal A}^{i,j})_{k,:}^{\top}]$ so that the constraint $\|\widehat{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]\|_{1}=1$} and $\widehat{\sigma}\left[({\cal A}^{i,j}_{k,l})\right]\geq 0$ is satisfied for the low-rank attention matrices in Fourier-KArAt. This algorithm is also kept as an optional sub-process for Algorithm 2.
Appendix C Addendum to Empirical Study
This section complements our empirical results in §5.
Table 5: Hyper-parameter settings for image classification and recognition experiments conducted in this work for vanilla ViTs and ViT+Fourier KArAts.
Input Size $224\times 224$
Crop Ratio $0.9$
Batch Size 128 for ImageNet-1K and 32 for CIFAR-10 & CIFAR-100
Optimizer AdamW
Optimizer Epsilon $1\times 10^{-6}$
Momentum $0.9$
Weight Decay $0.05$
Gradient Clip $1.0$
Learning Rate Schedule Cosine
Learning Rate $5\times 10^{-4}\times\frac{\text{Batch Size}}{512}$
Warmup LR $1\times 10^{-6}$
Min LR $1\times 10^{-5}$
Epochs 100
Decay Epochs 1
Warmup Epochs 5
Decay Rate $0.988$
Exponential Moving Average (EMA) True
EMA Decay $0.99992$
Random Resize & Crop Scale and Ratio $(0.08,1.0),(0.67,1.5)$
Random Flip Horizontal $0.5$; Vertical $0.0$
Color Jittering $0.4$
Auto-augmentation rand-m15-n2-mstd1.0-inc1
Mixup True
Cutmix True
Mixup, Cutmix Probability $0.5$, $0.5$
Mixup Mode Batch
Label Smoothing $0.1$
Patch Size 16
(a) Attention matrix $\mathcal{A}^{i,j}$} before softmax activation.
(b) Attention matrix $\sigma(\mathcal{A}^{i,j})$ after softmax activation.
Figure 5: Spectral analysis of the attention matrices before and after softmax shows that they are low-rank. For this experiment, we use all 3 heads in the last encoder block of ViT-Tiny on 5 randomly sampled images from the CIFAR-10 validation set. We plot all 15 singular vectors (each of 197 dimensions) where the singular values are arranged in non-increasing order.
C.1 Implementation Details
We implement the framework in Python using PyTorch [84] and use the training strategy of DEiT [14]. We train all models for $100$ epochs with the ADAM optimizer [85] with a base learning rate of $3.125\times 10^{-5}$ and batch-size $32$, except for the experiments on ImageNet-1K that use a learning rate of $1.25\times 10^{-4}$ with batch-size $128$. All experiments use a warm-up of $5$ epochs and a cosine scheduler with a weight decay of $5\times 10^{-2}$. The experiments were performed on two 80 GB NVIDIA H100 GPUs. The hyperparameter settings for experiments with ViTs are given in Table 5.
C.1.1 Baselines and Datasets
For our image classification benchmarking, we chose 5 popular vision Transformers—ViT-Tiny, ViT-Small [15, 16], ViT-Base [1], ConViT-Tiny [36], and Swin-Transformer-Tiny [37]. We incorporate Fourier-KArAt in them by replacing their softmax function with a learnable activation. We perform our benchmarking on CIFAR-10, CIFAR-100 [35] and ImageNet-1K [17] datasets. We use small-scale datasets, SVHN [38], Oxford Flowers-102 [39], and STL-10 [40] for understanding the transfer learning capability of the models. Additionally, we use ViT-Det [41] on the MS COCO [42] dataset for the object detection and segmentation task.
C.1.2 Small-Scale Datasets used
We use 3 small-scale datasets, SVHN [38], Oxford Flowers 102 [39], and STL-10 [40] for transfer learning task. SVHN is similar to MNIST [86], a 10-class digit classification dataset with small images like CIFAR-10 [35], divided into 73,257 training and 26,032 test images. The challenge in this dataset comes from the distracting elements surrounding the concerned class. The OxfordFlowers-102 [39] is a small yet challenging dataset with unbalanced training images (40-254 images) for the 102 classes. The STL-10 [40] is a CIFAR-like 10-class dataset with a larger image size, but only 500 training and 800 testing images per class.
C.1.3 Implementing Different Basis and Base Activation Functions in KArAt
[…]
$\textstyle{\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})=b({\cal A}^{i,j}_{k,q})+\sum_{m=1}^{G}a_{pqm}\cos{(m{\cal A}^{i,j}_{k,q})}+b_{pqm}\sin{(m{\cal A}^{i,j}_{k,q})}}.$ (5)
For the experiments in the main paper, in the general form, equation 5, we use $b(x)=0$ for all $x\in\mathbb{R}$, grid size, $G=3$, and the row dimension of the $\widehat{\Phi}^{i,j}$} is set to $r=12$, or otherwise noted. We encourage the reader to explore other configurations that differ from those previously outlined.
Different Basis Functions Used and Their Results. We incorporate a variety of widely used KAN activation functions into KArAt. By design, KArAt can utilize any basis function for activating the attention units. Therefore, in addition to the Fourier basis, we embed 5 different wavelet bases, and the Rational Function basis into KArAt. The function representations for these choices of bases are outlined in Table 6.
We use the safe Padé approximation unit (PAU) of order $(m,n)$ for rational function basis; KAT also uses the same basis function [3]. In contrast to PAUs, safe PAUs ensure stability and the prevention of poles, making it a clear choice for rational functions. We only consider safe PAUs of order $(5,4)$ in this paper; other order combinations can be used.
Being an integral part in signal analyses and image reconstruction, wavelet functions serve as a candid choice for approximating signals and feature extraction [48]. Having zero mean and finite energy drives wavelets to make fair and meaningful representations of signals in local regimes. Similar to the Fourier transform, the inverse continuous wavelet transform (CWT) resolves the original signal. We consider mother wavelets, mentioned in Wav-KAN [48], as basis functions. Unique to this choice, KArAt learns scale, $s$, translation, $\tau$, as measured in the CWT, and a learnable weight coefficient $w$. For any choice of wavelet base function, $\phi$, the input is first shifted by the scale and translation parameters and then activated by the mother wavelet.
Embedding different bases in other ViT variants with KArAt is straightforward. Although our original design in Algorithm 2 represents each matrix entry $\widehat{\phi}_{pq}$ using the Fourier basis, any functions outlined in Table 6 can be used. All other components in Algorithm 2 stay the same.
We implement rational and wavelet basis activation functions from Table 6 in ViT-Tiny+KArAt and ViT-Base+KARAT using the blockwise configuration and show their performance on the image classification task on CIFAR-10 and CIFAR-100; see the results in Table 8. Performance across model variants is not consistent with the choice of wavelet basis function used. ViT-Tiny+KArAt achieves its highest Top-1 accuracy with DOG (derivative of Gaussian) on CIFAR-10 and -100, whereas Rational (5,4) breaks this trend with ViT-Base+KArAt. Another observation shows that ViT-Base+KArAt performs poorly with both Meyer and Morlet functions, contrary to ViT-Tiny+KArAt. Hence, we cannot make a proper remark regarding the choice of basis.
Base Activation Functions Used and Their Results. We implement different base functions in ViT-Tiny+KArAt using the blockwise and universal configurations and show their performance on the image classification task on CIFAR-10; see the results in Table 8. For the majority of our experiments, including those in the main paper, we define and use ZeroModule and set the base activation function to be identically zero, or $b(x)=0$. Slightly lower computations within a single KAN layer help motivate us to use this base activation function. Overall, choosing between the activations SiLU and GELU for ViT-Tiny+KArAt with $G_{3}B,G_{3}U$ modes results in a slightly lower Top-1 accuracy compared to the results using the ZeroModule in the main paper (see Table 2).
Table 6: Different basis functions and their representations for $b(x)=0$. * For our experiments, the morlet central frequency hyperparameter is $\omega_{0}=5$, but other nonnegative values can be used. ** Meyer is defined to be $)(m\circ\nu)(\tilde{x})=m(\nu(\tilde{x}$)), where $m(t)=\mathbb{I}\left(-\infty,\frac{1}{2}\right](t)+\mathbb{I}{\left(\frac{1}{2},1\right)}(t)\cos\left(\frac{\pi}{2}\nu(2t-1)\right)$ and $\nu(x)=x^{4}(35-84x+70x^{2}-20x^{3})\mathbb{I}{[0,1]}(x)$.
Basis Function Representation, $\phi(x)$ Initialized Specifications
Fourier $\sum\limits_{k=1}^{G}\big(a_{k}\cos{(kx)}+b_{k}\sin{(kx)}\big)$ $a_{k},\ b_{k}\sim\mathcal{N}(0,1),\ G\text{ denotes the grid size}$
Rational $(m,n)$ $\frac{a_{0}+a_{1}x+\cdots+a_{m}x^{m}}{1+|b_{1}x+\cdots+b_{n}x^{n}|}$ $a_{i},\ b_{j}\sim\mathcal{N}(0,1),\ i=0,\ldots,m,\ j=1,\ldots,n$
Mexican Hat $\frac{2w}{\pi^{1/4}\sqrt{3}}\left(\tilde{x}^{2}-1\right)e^{-\frac{\tilde{x}^{2}}{2}},\ \tilde{x}=\frac{x-\tau}{s}$ $w\sim\mathcal{N}(0,1),\ \text{(Translation)~}\tau=0,\ \text{(Scale)~}s=1$
Morlet* $w\cos{\left(\omega_{0}\tilde{x}\right)}e^{-\frac{\tilde{x}^{2}}{2}},\ \tilde{x}=\frac{x-\tau}{s}$
DOG $w\frac{d}{dx}\left(e^{-\frac{\tilde{x}^{2}}{2}}\right),\ \tilde{x}=\frac{x-\tau}{s}$
Meyer** $w\sin{\left(\pi|\tilde{x}|\right)}\left(m\circ\nu\right)(\tilde{x}),\ \tilde{x}=\frac{x-\tau}{s}$
Shannon $w\text{sinc}\left(\frac{\tilde{x}}{\pi}\right)\omega(\tilde{x}),\ \tilde{x}=\frac{x-\tau}{s}$ $\omega(\tilde{x})\ \text{is the symmetric hamming window}$
Table 7: Blockwise and universal ViT-Tiny+KArAt and their Top-1 accuracies on CIFAR-10 using different base activation functions (SiLU, Identity, GELU).
  Model   Base Activation   Acc.@1
  on CIFAR-10
[…]
  SiLU   76.25
  GELU   76.48
[…]
  SiLU   75.25
  GELU   75.28
[…]
  SiLU   75.53
  GELU   75.18
[…]
  SiLU   75.12
  GELU   74.82
Table 8: Blockwise configuration ViT+KArAt and their Top-1 accuracies using different base functions on CIFAR-10 and CIFAR-100.
Model Basis Function Acc.@1
CIFAR-10 CIFAR-100
ViT-Tiny+KArAt Rational (5,4) 69.22 37.93
Mexican Hat 69.62 38.92
Morlet 70.84 37.88
DOG 73.89 44.23
Meyer 71.13 41.65
Shannon 71.41 39.40
ViT-Base+KArAt Rational (5,4) 72.59 41.52
Mexican Hat 72.58 44.44
Morlet 24.19 6.22
DOG 72.28 48.69
Meyer 28.83 6.67
Shannon 69.84 39.98
C.1.4 Deploying Fourier KArAt to other ViTs
We consider three modern ViT-based architectures, including ConViT [36] and Swin Transformer [37]. Incorporating the Fourier KArAt variants into these models is not straightforward. Below, we outline their implementation details.
ConViT [36] uses two different attention mechanisms, the gated positional self-attention (GPSA) and MHSA. All of the variants of ConViT vary the number of attention heads per encoder while keeping the number of encoders the same. The purpose of the ConViT is to combine a CNN and a ViT in one encoder block. With the first ten layers using GPSA and the last two layers using MHSA, the model learns a balanced structure between CNNs and ViTs. Our variant of ConViT replaces the MHSAs with Fourier KArAts. We implement the universal and blockwise attention modes on this model. We give the ConViT’s hyperparameter configuration details in Table 9.
Swin Transformers [37] has a variable number of tokens. The number of tokens considered in the attention operation varies based on the window size, and thus, we do not implement a universal attention mode due to incompatibility. We replace regular and shifted window MHSAs, W-MHSA and SW-MHSA, with regular and shifted Fourier KArAts. We give the Swin-Transformer’s hyperparameter configuration details in Table 10.
C.1.5 Implementing Fourier KArAt for Object Detection and Semantic Segmentation
Implementing Fourier KArAt for detection and segmentation is an involved task. Firstly, RCNN [87, 88, 89] frameworks employ an augmentation strategy to train on dynamic image sizes, which also dynamically change the number of tokens and thus, token interactions in the attention. However, KArAt needs to ensure fixed-length input and output; hence, the dynamic resize becomes incompatible. Secondly, vanilla ViT-Det [41] involves window partitioning in the self-attention, making the number of tokens variable in the attention matrices throughout the blocks of encoders. Thirdly, as this window mechanism changes the number of token interactions in the attention matrices, the universal mode of KArAt also becomes incompatible. To solve all these, we restrict the input image size to $224\times 224$, and use a window size of $14$ and a patch size of $16$. This inherently fixes the number of tokens to $196$, and the window size of $14$ ensures $196$ tokens in the windowed attention blocks. Thus, the number of tokens throughout the model remains fixed. We note that the restriction on augmentation and limiting the input size impacts the baseline performance of ViT-Det. The performance analysis is detailed in §5.1 and inference visualization is provided in Figure 8.
Table 9: Hyper-parameter settings for image classification conducted in this work for vanilla ConViT-Tiny and ConViT-Tiny+Fourier KArAt.
   Input Size    $224\times 224$
   Batch Size    128 for Imagenet-1K & 32 for CIFAR-10
   Optimizer    AdamW
   Optimizer Epsilon    $1\times 10^{-8}$
   Momentum    $0.9$
   Weight Decay    $0.05$
   Gradient Clip    $1.0$
   Learning Rate Schedule    Cosine
   Learning Rate    $5\times 10^{-4}\times\frac{\text{Batch Size}}{512}$
   Warmup LR    $1\times 10^{-6}$
   Min LR    $1\times 10^{-5}$
   Epochs    100
   Decay Epochs    30
   Warmup Epochs    5
   Decay Rate    $0.1$
   Exponential Moving Average (EMA)    False
   EMA Decay    $0.99996$
   Locality up to Layer    10
   Locality Strength    1
   Random Resize & Crop Scale and Ratio    $(0.08,1.0),(0.67,1.5)$
   Random Flip    Horizontal $0.5$; Vertical $0.0$
   Color Jittering    $0.4$
   Auto-augmentation    rand-m9-mstd0.5-inc1
   Mixup    True
   Cutmix    True
   Mixup, Cutmix Probability    $0.5$, $0.5$
   Mixup Mode    Batch
   Label Smoothing    $0.1$
Table 10: Hyper-parameter settings for image classification conducted in this work for vanilla Swin-Transformer-Tiny and Swin-Transformer-Tiny+Fourier KArAt.
   Input Size    $224\times 224$
   Crop Ratio    $0.9$
   Batch Size    128 for Imagenet-1K & 32 for CIFAR-10
   Optimizer    AdamW
   Optimizer Epsilon    $1\times 10^{-8}$
   Momentum    $0.9$
   Weight Decay    $0.05$
   Gradient Clip    $1.0$
   Learning Rate Schedule    Cosine
   Learning Rate    $5\times 10^{-4}\times\frac{\text{Batch Size}}{512}$
   Warmup LR    $1\times 10^{-6}$
   Min LR    $1\times 10^{-5}$
   Epochs    100
   Decay Epochs    30
   Patience Epochs    10
   Warmup Epochs    5
   Decay Rate    $0.1$
   Exponential Moving Average (EMA)    False
   EMA Decay    $0.99996$
   Fused Layer Norm    False
   Fused Window Process    False
   Window Size    7
   Random Resize & Crop Scale and Ratio    $(0.08,1.0),(0.67,1.5)$
   Random Flip    Horizontal $0.5$; Vertical $0.0$
   Color Jittering    $0.3$
   Auto-augmentation    rand-m9-mstd0.5-inc1
   Mixup    True
   Cutmix    True
   Mixup, Cutmix Probability    $0.5$, $0.5$
   Mixup Mode    Batch
   Label Smoothing    $0.1$
(a) ViT-Tiny
(b) ViT-Small
(c) ViT-Base
Figure 6: Training loss and test accuracy of vanilla ViTs and their Fourier KArAt versions on CIFAR-10, CIFAR-100, and ImageNet-1K datasets (left to right).
Table 11: Hyper-parameter settings for object detection and instance segmentation experiments.
Input Size $224\times 224$
Batch Size 32
Optimizer AdamW
Optimizer Epsilon $1\times 10^{-6}$
Momentum $0.9,0.999$
Weight Decay $0.1$
Learning Rate Schedule Warmup Scheduler
Learning Rate $1\times 10^{-4}$
Iterations 184375
Decay Rate $0.988$
Random Flip Horizontal $0.5$; Vertical $0.0$
Patch Size 16
Attention Window Size 14
Window Attention Block Indices $0,1,3,4,6,7,9,10$
Encoder Output Layer Index $11$
Pyramid Scale Factors $4.0,2.0,1.0,0.5$
Output Channels 256
Proposal Generator Input Layers (corresponding to feature pyramid) $p2,p3,p4,p5,p6$
Proposal Generator Input Sizes $32,64,128,256,512$
Proposal Generator Training Pre-NMS Top-K $12000$
Proposal Generator Evaluation Pre-NMS Top-K $6000$
Proposal Generator Training Post-NMS Top-K $2000$
Proposal Generator Evaluation Post-NMS Top-K $1000$
Proposal Generator NMS Threshold $0.7$
Region-of-Interest Heads IOU Threshold $0.5$
Region-of-Interest Score Threshold $0.05$
Region-of-Interest NMS Threshold $0.5$
C.1.6 Plotting Loss Landscapes
Following [67], we perform principal component analysis (PCA) of the parameter change over training progression to understand the major directions of parameter convergence. Considering the fully trained model as the minimum in the loss hyperplane and the two principal component directions as $X$ and $Y$ axes, we plot the loss values over the validation set of CIFAR-10 along the $Z$ axis for ViT-Tiny and -Base for the traditional MHSA and KArAt ($G_{3}B$ and $G_{1}B$, respectively).
C.2 Generalizability: Detection & Segmentation—Continued
We choose Mask RCNN [87]-based framework for this purpose and employ a feature pyramid network [90] based on ViT-Det [41]; it is non-trivial to implement Fourier KArAt, see §C.1.5. We trained ViT-Det [41] using ViT-Base backbone on the MS COCO [42] dataset for 50 epochs with two settings: (i) fine-tuning on the ImageNet-1K pre-trained weights on traditional softmax attention, and (ii) random initialization. Table 12 shows that for the fine-tuning on ImageNet pre-trained weights, Fourier-KArAt shows $\sim 4-13\%$ gap in average precision (AP) from its vanilla variant; their qualitative performance in Figure 8 is similar. Overall, for ViT-Det, Fourier-KArAt performs inferiorly to its conventional counterpart in detection and segmentation. Interestingly, Figure 9 shows, vanilla ViT-Det using softmax activation reaches its peak performance quickly (within 12 epochs), Fourier-KArAt delays in achieving it (within 30 and 45 epochs for $G_{1}B$ and $G_{1}U$, respectively), proving the incompatibility of softmax-based weights initialization. It is, however, still a good initialization for KArAt as the performance is better than random initialization.
Table 12: Fourier KArAt on object detection and instance segmentation tasks on the MS COCO [42] dataset. The header Box and Mask refer to detection and segmentation tasks, respectively.
Model Initialization Box Mask
AP AP50 AP75 AP AP50 AP75
ViT-Det-Base ViT-Base on ImageNet-1K 32.34 49.16 34.43 28.35 46.04 29.48
+ $G_{1}B$ 22.48 36.82 23.13 20.11 34.09 20.46
+ $G_{1}U$ 26.68 43.25 27.88 24.01 40.21 24.66
ViT-Det-Base Random Initialization 15.28 26.88 15.22 13.39 24.34 12.86
+ $G_{1}B$ 10.32 18.77 10.05 8.94 16.51 8.50
+ $G_{1}U$ 10.99 20.04 10.82 9.69 17.99 9.25
C.3 Computation Time, FLOPS, Memory Requirement, and Throughput of Fourier KArAt
The overall computation for Fourier KArAt variants is higher than their conventional softmax MHSA, and we have delineated it in Figure 7. Primarily, the Fourier KArAt variants have a longer training time. We show the training time comparison between the traditional MHSA and its Fourier KArAt versions for 100 epochs on CIFAR-10, CIFAR-100, and ImageNet-1K datasets for all the models (ViT-Tiny, -Small, and -Base) in Figures 7– 7, respectively. We only compare the best-performing Fourier KArAt models ($G_{1}B$ and $G_{1}U$ for ViT-Base, and $G_{3}B$ and $G_{3}U$ for ViT-Tiny & ViT-Small) with their traditional softmax MHSA counterparts. We also observe that universal mode $G_{n}U$ training times are consistently slightly less than the blockwise modes $G_{n}B$.
During the training, we monitored the GPU memory requirements, and as expected, Fourier KArAt variants utilize significantly more memory than traditional MHSA. In particular, the GPU memory requirements scale by $2.5-3\times$, compared to the traditional softmax MHSA.
We also compare the throughput during inference in Figure 7 and see slightly faster inference in universal mode than blockwise, except for ViT-Base. While there is a massive training time discrepancy between vanilla ViTs and Fourier KArAt ViTs, the inference speeds for Fourier KArAt variants are comparable to their vanilla counterparts. Although there is a minor difference in throughput between universal and blockwise modes during inference, theoretically, both variants for any model with the same grid size should have the same number of FLOPs.
(a) Training time on CIFAR-10 dataset.
(b) Training time on CIFAR-100 dataset.
(c) Training time on ImageNet-1K dataset.
(d) Throughput comparison during inference.
Figure 7: A detailed comparison of computing requirements.We compare the training times for 100 epochs with the hyperparameter settings given in Table 5 for all the datasets CIFAR-10, CIFAR-100, and ImageNet-1K. We also compare the throughputs of different models on ImageNet-1K; the throughput results will be similar for other datasets, as the input size is $224\times 224$.
Table 13: Parameter, computation, and memory requirement for Fourier-KArAt (with hidden dimension, $r=12$) compared to the traditional softmax attention. This Table particularly shows the individual computation required for the attention activation. The memory requirement shown is approximate and is based on averages of batches of $32$ images of resolution $224\times 224$. We note that changing $r$ will affect the performance and memory requirements. In our main paper, all the experiments were performed with $r=12$.
Model Parameters GFLOPs GPU Memory
Attention Activation Total Attention Activation Total
ViT-Base 0 85.81M 0.016 17.595 7.44 GB
+ $G_{1}B$ 1.70M 87.51M 0.268 17.847 17.36 GB
+ $G_{1}U$ 0.14M 85.95M 0.268 17.847 16.97 GB
ViT-Small 0 22.05M 0.008 4.614 4.15 GB
+ $G_{3}B$ 1.53M 23.58M 0.335 4.941 11.73 GB
+ $G_{3}U$ 0.13M 22.18M 0.335 4.941 11.21 GB
ViT-Tiny 0 5.53M 0.005 1.262 2.94 GB
+ $G_{3}B$ 0.76M 6.29M 0.168 1.425 7.48 GB
+ $G_{3}U$ 0.06M 5.59M 0.168 1.425 7.29 GB
(a) ViT-Det with softmax attention
(b) ViT-Det Fourier KArAt $G_{1}B$
(c) ViT-Det Fourier KArAt $G_{1}U$
Figure 8: Detection and segmentation tasks inference visualization using ViT-Det with ViT-Base as backbone for traditional MHSA and Fourier KArAt. For each sample, the ground-truth is given on the right side and the inference is on the left.
(a) Performance on Object Detection
(b) Performance on Instance Segmentation
Figure 9: Training Curves of ViT-Det and Fourier KArAt in object detection and instance segmentation tasks while training from ImageNet-1K pretrained ViT-Base with traditional MHSA weights initialization. In this particular task, the universal configuration ($G_{1}U$), performs strictly better than the blockwise configuration ($G_{1}B$).
C.4 Ablation Study with KAN and Fourier KArAt
In this section, we perform a detailed ablation study with different hyperparameter settings. Our first set of experiments shows why B-Splines are not a good basis for KAN for image classification tasks. The experiments with B-Splines solidify our argument for why we primarily use the Fourier basis in KArAt. After that, we perform rigorous ablation studies of Fourier KArAt on the hidden dimension, grid size effects, and many others.
C.4.1 Experiments with KANs Operated on B-Spline Basis
We test the performance of the B-spline KANs in the classification task on the CIFAR-10 [35], CIFAR-100 [35], MNIST [91, 86], and Fashion-MNIST [92] datasets. We performed extensive experiments to see if there are any benefits to choosing B-Splines for the basis functions in KAN layers. Table 14 shows the Top-1 accuracies from different variants of a Deep KAN. In this set of experiments, we closely followed earlier works [5, 6]. Hyperparameters involved in these experiments include the order of the B-spline $k$, grid size $G$, grid range [I,I][-I,I], layers, and width. Although the KANs with B-Spline basis yield high accuracies in the small-scale MNIST and Fashion-MNIST datasets, they fail to generalize over larger datasets (CIFAR-10 and -100).
Table 14: Experimental results on multi-layer KANs organized similarly to an MLP. These experiments involve B-Splines as the basis functions, as mentioned in Section 2.
(a) Experiments on CIFAR-10 and CIFAR-100.
Spline Order Grids Grid Range Layers Acc.@1
CIFAR-10 CIFAR-100
1 10 (-4, 4) (10, 10) 22 1
20 (-4, 4) (10, 10) 21 4
(-6, 6) (10, 10) 22 4
(-8, 8) (10, 10) 22 4
40 (-4, 4) (10, 10) 22 1
(-6, 6) (10, 10) 22 1
(-8, 8) (10, 10) 22 1
3 10 (-4, 4) (10, 10) 22 1
20 (-4, 4) (10, 10) 22 4
(-6, 6) (10, 10) 22 4
(-8, 8) (10, 10) 39 4
40 (-4, 4) (10, 10) 22 1
(-6, 6) (10, 10) 29 1
(-8, 8) (10, 10) 31 1
4 10 (-4, 4) (10, 10) 21 1
5 10 (-1, 1) (10, 10) 10 4
(-2, 2) (10, 10) 10 1
(-4, 4) (10, 10) 22 4
(-4, 4) (20, 20) 40 4
(-4, 4) (40, 40) 45 9
(-6, 6) (10, 10) 39 7
(-8, 8) (10, 10) 42 9
20 (-4, 4) (10, 10) 10 1
40 (-4, 4) (10, 10) 21 1
(b) Experiments on MNIST and Fashion-MNIST
Spline Order Grids Grid Range Layers Acc.@1
MNIST FMNIST
1 10 (-4, 4) (10, 10) 92 85
3 10 (-4, 4) (10, 10) 94 86
4 10 (-4, 4) (10, 10) 94 86
5 10 (-1, 1) (10, 10) 89 83
(-2, 2) (10, 10) 90 95
(-4, 4) (10, 10) 95 86
(-4, 4) (20, 40) 96 87
(-4, 4) (40, 40) 96 88
(-6, 6) (10, 10) 95 86
(-8, 8) (10, 10) 95 86
20 (-4, 4) (10, 10) 93 85
40 (-4, 4) (10, 10) 91 83
C.4.2 Ablation with the Hidden Dimension, $r$ in Fourier KArAt
While avoiding the computational overhead for computing $\Phi^{i,j}\in\mathbb{R}^{N\times N}$, we make use of the low-rank structure that the attention heads show (see Figure 5) by comparing different values of $r$. Particularly, we consider the values, $r=8,12,24$, on the Fourier KArAt variant of ViT-Base model, $\widehat{\Phi}^{i,j}\in\mathbb{R}^{r\times N}$; see Table 17 for results on CIFAR-10. In this ablation, we observe that changing the hidden dimension has a negligible impact on the model’s performance. This can be explained by the sudden drop in the singular values, as shown in Figure 13. As long as $r$ remains greater than the sudden drop index, the model should not be impacted by the changing of $r$ except for changes in computational requirement; a higher $r$ would incur a higher compute time as the size of the operator $\widehat{\Phi}$ scales with $r$.
Table 15: Ablation on grid size $G$ for ViT-Tiny, ViT-Small and ViT-Base. We find a particular grid size suitable for each of the models.
   Model    Acc.@1
   CIFAR-10    CIFAR-100
   ViT-Base    83.45    58.07
   + $G_{1}B$    81.81    55.92
   + $G_{1}U$    80.75    57.36
   + $G_{3}B$    80.09    56.01
   + $G_{3}U$    81.00    57.15
   + $G_{5}B$    79.80    54.83
   + $G_{5}U$    81.17    56.38
   + $G_{11}B$    50.47    42.02
   + $G_{11}U$    40.74    39.85
   ViT-Small    81.08    53.47
   + $G_{1}B$    79.00    53.07
   + $G_{1}U$    66.18    53.86
   + $G_{3}B$    79.78    54.11
   + $G_{3}U$    79.52    53.86
   + $G_{5}B$    78.64    53.42
   + $G_{5}U$    78.75    54.21
   + $G_{11}B$    77.39    52.62
   + $G_{11}U$    78.57    53.35
   ViT-Tiny    72.76    43.53
   + $G_{1}B$    75.75    45.77
   + $G_{1}U$    74.94    46.00
   + $G_{3}B$    76.69    46.29
   + $G_{3}U$    75.56    46.75
   + $G_{5}B$    75.85    –
   + $G_{5}U$    74.71    –
   + $G_{7}B$    75.11    –
   + $G_{7}U$    74.45    –
   + $G_{9}B$    74.85    –
   + $G_{9}U$    73.97    –
   + $G_{11}B$    74.52    –
   + $G_{11}U$    73.58    –
Table 16: Ablation on hidden dimension $r$ on CIFAR-10 with ViT-Base. Here, we compare the values of $r\in\{8,12,24\}$.
Model $r$ Acc.@1 on CIFAR-10
ViT-Base 58.07
+ $G_{3}B$ 24 80.54
+ $G_{3}U$ 24 80.81
+ $G_{5}B$ 24 77.99
+ $G_{5}U$ 24 80.52
+ $G_{3}B$ 12 80.09
+ $G_{3}U$ 12 81.00
+ $G_{5}B$ 12 79.80
+ $G_{5}U$ 12 81.17
+ $G_{3}B$ 8 80.76
+ $G_{3}U$ 8 80.40
+ $G_{5}B$ 8 79.79
+ $G_{5}U$ 8 80.83
Table 17: Comparing performance of Fourier KArAt for ViT-Tiny and ViT-Base with or without using $\ell_{1}$ projection in Algorithm 1. ViT-Tiny and ViT-Base use traditional softmax and do not require $\ell_{1}$ projection.
Model $\ell_{1}$ Projection Acc.@1
on CIFAR-10
ViT- Tiny 72.76
+ $G_{3}B$ 41.99
+ $G_{3}U$ 40.85
+ $G_{3}B$ 76.69
+ $G_{3}U$ 75.56
ViT-Base 83.45
+ $G_{3}B$ 47.44
+ $G_{3}U$ 46.11
+ $G_{3}B$ 80.09
+ $G_{3}U$ 81.00
Table 18: Complete compute requirement of ViT-Tiny+Fourier KArAt with extremely low hidden dimension training.
Model $r$ Acc.@1 Parameters GFLOPs GPU Memory
CIFAR-10 CIFAR-100
ViT-Tiny 72.76 43.53 5.53M 1.262 2.94GB
+ $G_{3}B$ 12 76.69 46.29 6.29M 1.425 7.48GB
4 75.90 45.76 5.80M 1.313 4.68GB
2 74.39 44.36 5.56M 1.285 4.01GB
+ $G_{3}U$ 12 75.56 46.75 5.59M 1.425 7.29GB
4 73.13 44.98 5.57M 1.313 5.08GB
2 71.56 42.43 5.55M 1.285 4.41GB
ViT-Base 83.45 58.07 85.81M 17.595 7.44GB
+ $G_{1}B$ 12 81.81 55.92 87.51M 17.847 17.36GB
4 81.20 55.29 86.41M 17.668 12.05GB
2 81.55 56.49 86.13M 17.623 9.48GB
+ $G_{1}U$ 12 80.75 57.36 85.95M 17.847 16.97GB
4 79.69 55.70 85.89M 17.668 12.11GB
2 80.36 54.04 85.87M 17.623 9.92GB
C.4.3 Enforcing an Extreme Low Rank Structure in Fourier KArAt
[…]
C.4.4 Ablation on the Impact of grid size, $G$ on Fourier KArAt
KANs are highly dependent on certain hyperparameters, and Fourier-KArAt has only one hyperparameter to tune the performance — grid size $G$. Thus, we perform extensive experiments involving grid size $G$ and present them in Table 17. We observe that each of the particular ViT models, in conjunction with particular Fourier-KArAt variants, has a typical $G$ value that brings out its best performance, and there is no universal value of $G$ to follow. When performing validation with ViT-Base+ variants on CIFAR-10 and CIFAR-100, the accuracy drops as the grid size passes a size after 5. However, this behavior is not persistent with the ViT-Small/Tiny+ variants; see Table 17333We did not perform gridsize 5, 7, 9, and 11 experiments with ViT-Tiny+KArAt as it was already outperforming the base ViT-Tiny with gridsize 1 and 3, and the experiments are extensively resource intensive..
Figure 10: Optimization path for ViT-Tiny and ViT-Base (the smallest and the largest model) along the two largest principal component directions of the successive change of model parameters. We show the loss contours along with the trajectory of the optimizers. Perturbed contours indicate corresponding non-smooth, spiky loss surfaces.
Figure 11: Vit-Tiny attention map characterization. Original image for inference (the center), the attention maps (top row), and contributing image regions (bottom row) for all three heads in ViT-Tiny: traditional MHSA (left) and Fourier KArAt $G_{3}B$ (right). The traditional MHSA sporadically focuses on fine-grained features of the primary object in the image. In contrast, the learnable attention in Fourier KArAt identifies the primary object features present significantly across all heads.
Table 19: Ablation study on the operator variants in Fourier KArAt. Note that $\widehat{\Phi}(W{\cal A})$} and $W\widehat{\Phi}({\cal A})$} refer to order of using linear projector $W$ and operator $\widehat{\Phi}$, and $\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ refers to two consecutive learnable operators $\widehat{\Phi}_{[1]}$ and $\widehat{\Phi}_{[2]}$ applied to $\cal A$}.
Model Acc.@1 on CIFAR-10
$\widehat{\Phi}(W{\cal A})$ $W\widehat{\Phi}({\cal A})$ $\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$
ViT-Base
+ $G_{11}B$ 41.02 50.47 30.14
+ $G_{11}U$ 47.74
ViT-Tiny
+ $G_{11}B$ 67.04 75.41 33.09
+ $G_{11}U$ 70.98 73.67 36.98
Table 20: Performance using softmax and variants of Fourier KArAt in separate heads.
Model Low Rank Configurations No. of Distinct Heads Acc.@1 on
(for Fourier KArAt) softmax Fourier KArAt CIFAR-10
ViT-Base 12 0 83.45
ViT-Base + $G_{11}B$ $\hat{\sigma}=\widehat{\Phi}(W{\cal A})$ 0 12 41.02
$\hat{\sigma}=\widehat{\Phi}(W{\cal A})$ 6 6 47.60
$\hat{\sigma}=W\widehat{\Phi}({\cal A})$ 0 12 50.47
$\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 0 12 30.14
$\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 6 6 30.24
ViT-Tiny 3 0 72.76
ViT-Tiny + $G_{11}B$ $\hat{\sigma}=\widehat{\Phi}_{[1]}(\widehat{\Phi}_{[2]}({\cal A})$ 0 3 33.09
$\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 1 2 32.51
$\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 2 1 34.73
ViT-Tiny + $G_{11}U$ $\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 0 3 36.98
$\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 1 2 33.12
$\hat{\sigma}=\widehat{\Phi}_{[2]}(\widehat{\Phi}_{[1]}({\cal A}))$ 2 1 70.57
C.4.5 Does Fourier KArAt Require $\ell_{1}$-Projection?
To ensure that each row vector of the learned attention matrix lies on a probability simplex, we project it onto the $\ell_{1}$-unit ball. We use Algorithm 1 to project learned attention vectors to the probability simplex and compare Top-1 accuracies to the baseline model. We note that using $\ell_{1}$ projection does not substantially increase the training time. However, from Table 17, we observe that incorporating this algorithm in the Fourier-KArAt does not improve its performance; instead, the performance significantly degrades. Algorithm 1 can be used with the choice of any basis function. However, we cannot comment on the role of $\ell_{1}$-projection in KArAt’s performance with basis functions other than Fourier.
C.4.6 More Discussion on Blockwise and Universal Configuration
Although the primary inspiration behind the universal configuration comes from the shared basis functions in KAT [3], it aligns with the fact that softmax is a fixed function conventionally being used across all heads of all encoder layers in the ViT architectures. This raises the question of whether the token-to-token interaction in the self-attention can be modeled using a single universal function, or each instance of attention operations needs dynamic modeling using learnable functions. To investigate further, we compared the learned coefficients $\{a_{pqm}$, $b_{pqm}\}$ from Equation 2 and the weight matrix $W$ of the linear projector following the learnable basis, for each head across all the encoder layers in ViT-Tiny+Fourier KArAt. Thus, we obtain 36 sets (3 heads for each of the 12 layers) of such coefficients and weights from the model with blockwise configuration $G_{3}B$, and 3 sets from the model with universal configuration $G_{3}U$. Although mostly centered at zero, the distribution of each of these sets of parameters varies significantly from the others. The difference is mostly visible in the presence of long tails and the length of the tails. This characteristic is present not only in the blockwise configuration, but the three heads of the universal configuration also show a similar pattern.
C.5 Alternate Variants of Fourier KArAt
In this section, we experiment with different attention variants on Fourier KArAt.
C.5.1 Alternative Approaches to the Lower Rank Attention Structures
[…]
Primarily, we started with the full-rank operator $\Phi\in\mathbb{R}^{N\times N}$ and found its computations to be prohibitively expensive, regardless of the choice of the basis. Next, we conduct an ablation study to see which operator configuration works better; see Table 20. Our experiments show that the operator configuration (ii) demonstrates an inferior performance. We postulate that by down-projecting the $N$-dimensional attention row vector to a smaller dimension, $r$ loses adequate token-to-token interaction, and this fails to capture dependencies from significant attention units within a row. After that, from this limited information, learnable activation cannot significantly help the model’s performance. On the other hand, (iii) also fails to perform due to its inability to model the attention well, despite having refined information, from the intermediate $r$-dimensional subspace. Overall, apart from configuration (i), the performances of the other configurations were inconsistent over multiple experiments. Also, they lag in training stability, particularly configuration (iii). Considering these observations, we proceeded to find the best possible grid size $G$ only with the configuration (i).
C.5.2 Fourier KArAt + softmax Attention —A Hybrid version
We consider another variant where we mix the learnable and pre-defined activation in each head. E.g., it is possible to have 2 of the 3 attention heads in each encoder block in ViT-Tiny activated by softmax and the third activated by the Fourier KArAt. With the idea of incorporating KAN to replace softmax activation, we are curious to see if the ViT model performs better with a hybrid mode of activation; see Table 20 for results.
C.5.3 More Combinations
We have also experimented with more configurations of Fourier KArAt in various permutations of the strategies mentioned above and have not found any significant combination in terms of performance. We also carefully design a particular training strategy where the attention operators ($\widehat{\Phi}$, $W$, and $\Phi$} wherever applicable) of Fourier KArAt are trained with a separate learning rate to alleviate the problem of the mismatch of the data passing through these learnable layers, as they consider each row of an attention matrix $\cal A$} as an input. The models have a gradient and loss explosion during training.
C.6 Distribution of the Weights
[74] considered an invariant measure perspective to describe the training loss stabilization of the neural networks. We adopt this idea to study the distribution of the weights of the smallest and largest models of the ViTs, Tiny and Base, and their Fourier KArAt variants. Figure 12 shows the distribution of the weights of these models during different training phases—The evolution of the weights’ distributions for respective models remains invariant. Based on this observation, from [74], we can guarantee the convergence of loss values of all models; also, see Figure 6. However, as mentioned in [74], this perspective cannot comment on the generalization capacity and structural differences between different neural networks.
(a) ViT-Tiny
(b) ViT-Tiny+Fourier KArAt
(c) ViT-Base
(d) ViT-Base+Fourier KArAt
Figure 12: Weight distribution of ViT-Tiny and -Base with traditional MHSA and Fourier KArAt. The columns (left to right) represent weights at initialization, epoch 50, and epoch 100.
C.7 Spectral Analysis of Attention
Although we cannot comment on the generalizability by studying the distribution of the weights, from Table 2, we realize all the KArAt variants have more parameters than their vanilla counterparts. Therefore, it would be interesting to see how their attention matrices behave. As discussed in §4, the attention matrices in traditional MHSA have a low-rank structure. Following that study, we verified the apparent low-rank structure that Fourier KArAt’s learned attention matrices possess.
We use all 3 heads in the last encoder block of ViT-Tiny on 5 randomly sampled images from the CIFAR-10 validation set. There are a total of 15 singular vectors (each of 197 dimensions) for any attention matrix of shape $197\times 197$, where the singular values are arranged in non-increasing order. Let $\sigma_{i}$ be the $i^{\rm th}$ singular value across all heads and samples (it is permutation invariant). For each $i\in[197],$ we plot $\ln(\sigma_{i}^{\min}),\ln(\bar{\sigma_{i}}),\ln(\sigma_{i}^{\max})]$, where $\bar{\sigma_{i}}$ is the average of $i^{\rm th}$-indexed singular value across all samples and heads.
To investigate the inherent low-rank structure of attention, we plot the natural logarithm of singular values of attention matrices, $\sigma_{i}$, before and after attention activation in Figure 13. From Figure 13, we observe that the traditional softmax attention and our learnable Fourier KArAt have almost similar low-rank structures. However, Figure 13 shows that the traditional MHSA has significantly larger singular values than its KArAt variant, $G_{3}B$. It can also be noticed that before the activation, both traditional MHSA and Fourier KArAt feature a sharp drop in singular values between the 50th and 75th indices. This sharp drop in singular values vanishes in softmax attention, indicating a normalization. However, due to the hidden dimension $r$, Fourier KArAt enforces a much lower rank than the traditional MHSA.
Additionally, we observe from the weight distributions in Figure 12 that Fourier KArAt variants have more entries close to zero than their ViT counterparts. This, along with the low-rankness, can inspire low-rank + sparse training strategies of Fourier KArAT.
(a) Before attention activation
(b) After attention activation
Figure 13: Spectral decomposition of the attention matrix for ViT-Tiny on CIFAR-10 dataset with traditional softmax attention and our learnable Fourier KArAt. The traditional softmax attention and our learnable Fourier KArAt have almost similar low-rank structure, before activation functions are used.